### v1

# CLEAR: Character Unlearning in Textual and Visual Modalities

Abstract
Machine Unlearning (MU) is critical for enhancing privacy and security in deep learning models, particularly in large multimodal language models (MLLMs), by removing specific private or hazardous information. While MU has made significant progress in textual and visual modalities, multimodal unlearning (MMU) remains significantly underexplored, partially due to the absence of a suitable open-source benchmark. To address this, we introduce CLEAR, a new benchmark designed to evaluate MMU methods. CLEAR contains 200 fictitious individuals and 3,700 images linked with corresponding question-answer pairs, enabling a thorough evaluation across modalities. We assess 10 MU methods, adapting them for MMU, and highlight new challenges specific to multimodal forgetting. We also demonstrate that simple $\ell_{1}$ regularization on LoRA weights significantly mitigates catastrophic forgetting, preserving model performance on retained data. The dataset is available at https://huggingface.co/datasets/therem/CLEAR.
CLEAR: Character Unlearning in Textual and Visual Modalities
Alexey Dontsov1,6, Dmitrii Korzh1,3, Alexey Zhavoronkin2,4, Boris Mikheev3, Denis Bobkov1,6, Aibek Alanov1,6, Oleg Y. Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,6 1AIRI 2MIPT 3Skoltech 4Sber 5University of Sharjah 6HSE University Correspondence: dontsov@airi.net; tutubalina@airi.net
1 Introduction
Figure 1: The overview of our dataset.
Figure 2: Summary of our dataset. We generate 200 persons and use multimodal unlearning to forget the part of them. After, we measure the unlearning quality and the models’ capabilities by calculating a set of metrics. Then, we create a leaderboard of unlearning methods based on these metrics.
Large Language Models (LLMs) Touvron et al. (2023); Jiang et al. (2023) are trained on vast corpora of data that contain private, unethical, or unwanted information, leading to growing concerns. Machine unlearning (MU) methods have been developed to remove such unwanted data without expensive retraining from scratch. For instance, MU has been applied for the LLMs to mitigate issues related to toxicity Lu et al. (2022), copyright and privacy concerns Jang et al. (2022); Eldan and Russinovich (2023); Wu et al. (2023) and fairness Yu et al. (2023). Additionally, such topics as model editing Ilharco et al. (2022); Zhang et al. (2023), prevention of hallucinations Yao et al. (2023), and sensitive knowledge exposure Barrett et al. (2023) have also motivated the development of MU techniques.
There are various unlearning techniques suitable solely for LLMs Yao et al. (2024b, a); Xing et al. (2024); Zhang et al. (2024) or for vision models Li et al. (2024a); Chen and Yang (2023); Tarun et al. (2021). However, multimodal LLMs (MLLMs) Liu et al. (2023), specifically visual LLMs (VLLMs), raise new challenges. The unlearning of such multimodal models (MMU) remains largely unexplored, primarily due to the lack of open-source benchmarks. Moreover, current MU benchmarks Maini et al. (2024); Shi et al. (2024); Yao et al. (2024a); Li et al. (2024b) are focused on single modalities, and, to our knowledge, no open benchmarks designed explicitly for evaluating unlearning in multimodal models exist at the time of submission.
To address this gap, we propose CLEAR, a new benchmark for textual-visual MMU named, focusing on person unlearning, which aligns with the right-to-be-forgotten concept. The dataset is synthetic to ensure control over the data the model learns, preventing object leakage during training. We generated consistent images through a comprehensive strategy and linked them to the corresponding author-related questions from the large-scale textual unlearning benchmark TOFU Maini et al. (2024). The proposed dataset contains 200 fictitious authors, 3,770 visual question-answer pairs, and 4,000 textual question-answer pairs, enabling a thorough evaluation of the single and multi-modal unlearning techniques. We also propose a benchmark to assess MU and MMU methods, evaluating 10 techniques, including the current state of the art.
We evaluate the unlearning methods in textual, visual, and multimodal setups. First, we finetune the model on our entire dataset. Next, we take a small predefined subset of 20 authors, referred to as the forget set, while the remaining data forms the retain set. We then apply the unlearning procedure, resulting in a new model that no longer "remembers" individuals from the forget set but retains knowledge of those in the retain set. To ensure the model’s capabilities are not compromised during unlearning, we assess its performance on real-world tasks, such as celebrity face recognition and general domain visual question-answering (VQA). Overall, we work with four sets: Forget, Retain, Real Faces, and Real World. Figure 1 illustrates an example from each set.
We evaluate existing unlearning methods separately in textual and vision modalities and then combine them within an MLLM. We create a leaderboard for each domain, highlighting that multimodal unlearning poses new challenges.
Finally, we demonstrate that applying $\ell_{1}$ regularization to the LoRA adapter during the unlearning process significantly improves performance, helping to prevent catastrophic forgetting of the retain set information. Overall, our contributions can be summarized as follows:
  • We propose a novel benchmark, CLEAR, for evaluating machine unlearning in multimodal (textual-visual) setups. To the best of our knowledge, this is the first publicly available MMU benchmark.
  • We comprehensively evaluate existing unlearning methods across separate and combined domains. We construct leaderboards for these three domains and show that state-of-the-art unlearning algorithms struggle in multimodal setups, highlighting the need for new approaches.
  • We demonstrate that the $\ell_{1}$ weight regularization on the LoRA adapter helps to improve unlearning quality by significantly preventing catastrophic forgetting.
2 Related Work
2.1 MU Methods and Textual Benchmarks
MU methods Cao and Yang (2015); Dwork et al. (2014); Kurmanji et al. (2024); Neel et al. (2021); Sekhari et al. (2021) remove the impact of certain data instances from a trained model without requiring full retraining. The goal is to obtain a model that behaves like the forget data was never part of the training set. MU can be formalized in two main ways. In a setting where the unlearned model must produce identical outputs to a model trained without the forget data. In an inexact unlearning setting, the main objective is to obtain any model that doesn’t contain knowledge from the forget set, but with no restrictions and guarantees on the model’s response to inputs from the retain set.
There are several standard textual unlearning benchmarks. TOFU Maini et al. (2024) is a benchmark for textual LLM unlearning, featuring 200 fictitious author profiles, each defined by attributes such as name, birthplace, parent’s names and occupation, written books, etc. Totally the dataset has 4,000 question-answer pairs (20 per author). Predefined 10/90, 5/95, and 1/99 splits are used for forget/retain sets. Models are finetuned on the entire dataset, and unlearning is applied to the forget pairs. WMDP Li et al. (2024b) consists of 3,668 multiple-choice questions, evaluates hazardous knowledge of LLMs, and serves as a benchmark for unlearning such dangerous information. In (Yao et al., 2024a), the authors explore seven textual unlearning techniques. The best approach was a combination of gradient ascent with gradient descent on the forget and retain sets, respectively, to maintain the unlearn/retain quality trade-off. They evaluated their methods on three benchmark datasets from different domains: 500 arXiv papers, 2000 GitHub files, and 100 books covering academic texts, code, and literary works. Unfortunately, these three benchmarks cannot be applied to the MMU evaluation.
Figure 3: Examples of generated images showcasing a distinct individual from our dataset.
2.2 MMU Methods and Benchmarks
MLLMs Liu et al. (2023) typically consist of three core components: a modality encoder that translates raw input into feature embeddings, a modality projection layer aligning these features within the language space, and a pre-trained language model synthesizing the final output.
Nonetheless, MMU research is still in its early stages. Cheng and Amiri (2023) proposed the MultiDelete method, which focuses on separating cross-modal embeddings for the forget set while preserving unimodal embeddings for the retain set. Unfortunately, this approach is suitable only for the encoder-decoder architecture and can not be directly transferred to decoder-only LLMs.
EFUF Xing et al. (2024) mitigates hallucinations in MLLMs using unlearning. It measures the similarity between generated captions and image content with CLIP model Radford et al. (2021) to automatically detect hallucinated (negative) and non-hallucinated (positive) examples based on the calibrated on the MSCOCO dataset Lin et al. (2014) thresholds, eliminating manual labeling. The unlearning process applies three loss functions: negative loss to forget hallucinations, positive loss to reinforce correct representations, and sentence loss to maintain fluency. However, their benchmark is not open-sourced.
Single Image Unlearning (SIU) Li et al. (2024a) focuses on unlearning visual concepts in MLLMs while preserving textual knowledge and introduces MMUBench with five evaluation metrics. The benchmark covers 20 concepts with at least 50 images for the concept, including real-world figures and cartoon characters. For each concept, one image is selected as the forget subset, paired with various prompts, while the remaining images form the retain subset. However, SIU’s use of a single image raises scalability concerns for complex concepts, and their unlearning of VLLM is limited to the visual domain. Moreover, MMUBench is also not open-sourced.
Chakraborty et al. (2024) explores unlearning harmful content in VLLMs, demonstrating that unlearning in the textual domain alone can match the performance of text-image unlearning while using fewer resources. Their method combines harmful loss and KL divergence between unlearned and retrained models. The approach was tested on six datasets, including PKU-SafeRLHF Ji et al. (2024) and three vision-text attack datasets Shayegani et al. (2023); Luo et al. (2024); Gong et al. (2023) for harmful content prompts, and Truthful-QA and VQA-v2 Lin et al. (2021); Goyal et al. (2017) to ensure benign task performance remained intact. However, their focus on safety alignment may limit applicability for general unlearning, such as biometric privacy. While they claim that textual unlearning is sufficient for MMU, our findings show this is not true for all methods. Additionally, their approach lacks exact unlearning evaluation.
3 Methodology
Unlearning can be conceptualized in two main ways: the objective of the Strict Unlearning is to achieve a model that behaves identically to one trained exclusively on the retain set, ensuring that no knowledge from the forget set is present; the objective of the Inexact Unlearning is to produce any model that no longer contains information from the forget set. However, this approach cannot guarantee how the model will respond to inputs related to the forget set. These objectives require distinct methods, loss functions, and evaluation metrics to assess unlearning quality.
Let $f_{\theta}$ denote the original model with its parameters $\theta$. Source model $f_{\theta}$ is trained on train dataset $D$, and given the unlearning objective, we want to make our model forget a subset of the source dataset $D$, called forget set $D_{F}$. The remaining part of the training dataset is called retain set, and we aim to preserve the model’s performance on this data subset $D_{R}:=D\setminus D_{F}$. Additionally, we utilize a holdout set $D_{H}$ to establish a reference for the model’s desired behavior on $D_{F}$ after the unlearning process. The model’s training did not include this set, ensuring that $D_{H}\cap D=\emptyset$. In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behavior on data that was not involved in the training process. Such forgetting procedure is performed by updating the model $f_{\theta}$ with a particular unlearning method, which results in a new unlearned model $f_{\hat{\theta}}$ with parameters $\hat{\theta}$. For the evaluation, we can also train separately a gold model $g_{\omega},$ which is trained only on the $D_{R}$.
So the objective is to obtain the unlearned model $f_{\hat{\theta}}$ which unlearns the forget set $D_{F}$ while preserving the performance on retain set $D_{R}$ as good as the source model $f_{\theta}$. It can be done by optimizing (minimizing or maximizing) the specific criterion with the model parameters $\theta$ and resulting obtained parameters $\hat{\theta}$ will be the parameters of desired unlearned model $f_{\hat{\theta}}$. For example, one can consider the gradient difference MU approach for the optimization, aimed at increasing forget loss and maintaining retain performance:
$\tilde{L}=-\sum_{x_{i}\in D_{f}}L(x_{i},y_{i},\theta)+\lambda\sum_{x_{j}\in D_{R}}L(x_{j},y_{j},\theta)$ (1)
$\theta\mapsto\theta-\alpha\nabla_{\theta}\tilde{L},$ (2)
where $\lambda$ – forget-retain trade-off hyper-parameter, $\alpha$ – learning rate, $L$ is a loss function, for example, negative-log-likelihood. $x$ is the input – text, image, or both of them in the case of VLLM.
In this work, we explore a subset of unlearning methods, including Retain Finetune, Gradient Ascend (GA), Gradient Difference (GD) Liu et al. (2022), SCRUB Kurmanji et al. (2023), DPO Rafailov et al. (2023), NPO Zhang et al. (2024), LLMU Yao et al. (2024b) , IDK Maini et al. (2024), RMU Li et al. (2024b), and KL Golatkar et al. (2020), which are detailed in Appendix A.We selected these methods based on their ease of adaptation to new modalities, requiring only changes in input data (text, images, or both) while preserving their core functionality. In essence, these methods involve variations of hard negative training on the forget set combined with fine-tuning on the retain set, often with additional constraints such as Cross-Entropy or KL divergence to align the model’s outputs on the retain set with the original model.
4 CLEAR
The MU (and consequently MMU) benchmark should ideally avoid running unlearning on well-known information that could be obtained from external sources such as books, games, movies, etc. This is essential for a more reliable evaluation of the model’s performance on retain and forget. To meet this requirement, we chose to extend the TOFU dataset due to its ease of use, flexibility for adopting to new modalities (such as adding face images or personal voices), and its strong connection to privacy concerns, making it ideal for testing unlearning in sensitive contexts.
4.1 Dataset Generation Process
Firstly, for each of the 200 authors from the TOFU dataset, we extract their name, age, and ethnicity based on the knowledge provided in the original dataset. Also, we generate a pool of 2000 faces using StyleGAN2 Karras et al. (2020) - an established generative model for face synthesis. Each face is scored with pre-trained CNNs to get the age, gender, and ethnicity of the face. Then, we manually select the most suitable face that matches these attributes for each author. During this phase, we discovered that the age distribution of the authors was highly shifted towards the older ages, so we needed to eliminate this gap. To do this, we used the image editing framework proposed in Bobkov et al. (2024) to shift the visual attributes of the faces to make them older. The distribution of the characteristics of the faces and authors is shown in Figure 4. After matching each author to a face, we used the diffusion model Li et al. (2024c) of personalized generation to synthesize images with a given face and corresponding to a given prompt. In detail, the face generation and collection process is described in Appendix B.
Figure 4: Distributions of the attributes of the author’s faces. We show that CLEAR is balanced and representative regarding age, gender, and ethnicity.
The diffusion model needs a textual prompt for image generation besides the face. We ask GPT-4 to generate images from a textual question and an answer about an author. We generate 8 images for each prompt, evaluate them using an ensemble of fake-detection models, and select the most realistic. Additionally, GPT-4o generates captions for each image and visual prompt pair, which are then included in the dataset. However, due to restrictions caused by GPT guard breaks and the identification of several bugs in the TOFU dataset (such as a nameless author), the final dataset includes fewer images than text pairs (3,770 compared to 4,000). We also incorporated two additional data splits containing real-world face photos and natural images to maintain the model’s visual capabilities during unlearning.
4.2 Splits
Ultimately, we utilize the following four splits (sets) to evaluate unlearning:
Forget. Following methodology from Maini et al. (2024), $D_{F}$ is made from data of 2, 10, and 20 persons (1%, 5% and 10% correspondingly) of the full set $D$, consisting of 200 authors. This $D_{F}$ is expected to be unlearned by the model.
Retain. Retain set $D_{R}$ is made from all the other data from the complete set $D$, not included in the $D_{F}$. The model should continue to work well on this subset and preserve its performance as much as possible.
Real Faces. To ensure the model retains knowledge of related concepts, such as faces, which are not present in the finetuning dataset, we evaluate it using a set of real-world faces. Specifically, we use the MillionCelebs dataset Zhang et al. (2020), which consists of celebrity face-name pairs. We intersect this dataset with the most recognized celebrities from any year on the Forbes Celebrity 100 list to increase the likelihood that the model has seen these faces during pre-training. This results in a final set of 150 face-name pairs.
Real World. To ensure that the model’s overall visual capabilities remain intact throughout the unlearning process, we evaluate its performance on the Visual Question Answering (VQA) task using samples from x.ai (2024).
Figure 1 presents a random sample from each of these splits.
4.3 Evaluation Metrics
To comprehensively evaluate unlearning performance on textual, visual, and textual-visual domains, we measure the MU and MMU performance in terms of the following metrics:
ROUGE-L We calculate the ROUGE-L Lin (2004) score between the model’s predictions and ground truth answers. This metric measures how much the model remembers in exact formulations. However, the generation of models does not always represent the inner knowledge of one’s. That is where the next metric is used.
Probability Score One way of exposing elicit knowledge from a model is through its logits, which are assigned to some factual tokens. We define the conditional probability $p(y|x)^{\frac{1}{|y|}}$ for input $x$ and answer $y$ (power $\frac{1}{|y|}$ corresponds to normalizing for length). Each input question $x$ is considered as a multiple choice question with possible answers $y_{1},...,y_{n}$, and then, assuming $y_{1}$ is the correct answer, desired probability score is computed as $\frac{p(y_{1}|x)}{\sum\limits_{i=1}^{n}p(y_{i}|x)}$. It will be bounded between 0 and 1. A lower probability score indicates that the model is less confident about generating content.
Truth Ratio quantifies the alignment between predictions and the ground truth by comparing the probability of a paraphrased correct answer against the averaged probabilities of several similarly formatted incorrect answers, providing insight into the effectiveness of the unlearning algorithm in removing specific information while maintaining overall accuracy. Assume that $\hat{y}$ denotes a paraphrased version of the answer $y$ for input $x$ and $Y^{\prime}$ is the set of 5 perturbations of the answer $y$. Then desired truth ratio $R$ is calculated as:
$R=\frac{\frac{1}{|Y^{\prime}|}\sum\limits_{y^{\prime}\in Y^{\prime}}p(y^{\prime}|x)^{\frac{1}{|y^{\prime}|}}}{p(\hat{y}|x)^{\frac{1}{|\hat{y}|}}}.$ (3)
This ratio is normalized and rescaled between 0 and 1, with higher values indicating improved unlearning.
Forget Quality. Measuring the quality of forgetting in MU presents significant challenges. The objective is to create a model that cannot be distinguished from one trained solely on retain. The established Hayes et al. (2024) way of measuring the unlearning quality is calculating the U-LIRA score. However, it requires training at least 128 model copies, which is computationally expensive for LLM. A feasible method for achieving this is proposed. We calculate a statistical test on the outputs of two models: our unlearned model and the gold model. The Truth Ratio metric is considered as output for its effectiveness in informativeness. To assess this metric, the Kolmogorov-Smirnov test (KS-Test) is employed to compare the distributions of Truth Ratios from both models. A high p-value from this test suggests effective forgetting, while a low p-value indicates potential privacy leakage and poor unlearning. We call this p-value the Forget Quality of the unlearning method.
Also, we define Real, Retain, and Forget metrics as a harmonical mean of ROUGE, Real Probability score, and Truth Ratio.
M Method Real Metric $\uparrow$ Retain Metric $\uparrow$ Forget Metric $\downarrow$ Log Forget Quality $\uparrow$
LLama2-7B Retain FT 0.50 0.26 0.42 -4.92
LLMU 0.38 0.03 0.01 -2.31
KL 0.24 0.00 0.00 -18.22
GA 0.25 0.00 0.00 -17.22
GD 0.61 0.13 0.01 -48.59
IDK 0.46 0.26 0.24 -4.92
DPO 0.50 0.26 0.42 -4.92
SCRUB 0.50 0.26 0.42 -4.92
RMU 0.51 0.26 0.59 -42.86
NPO 0.50 0.28 0.62 -44.46
Mistral-7B Retain FT 0.67 0.34 0.47 -3.87
LLMU 0.65 0.30 0.39 -6.69
KL 0.28 0.00 0.00 -50.30
GA 0.26 0.00 0.00 -36.06
GD 0.60 0.01 0.00 -51.16
IDK 0.63 0.32 0.45 -2.72
DPO 0.67 0.33 0.47 -3.63
SCRUB 0.66 0.33 0.47 -3.39
RMU 0.09 0.00 0.00 -123.22
NPO 0.67 0.33 0.47 -3.16
Table 1: Unlearning methods on textual domain only. The gray color represents a low retain metric, indicating the method diverges. Hence, we do not consider them.
Method Forget Acc. $\downarrow$ Holdout Acc. $\uparrow$ Retain Acc. $\uparrow$ U-LIRA $\downarrow$ U-MIA $\downarrow$
Original 100.00 18.50 100.00 1.00 0.96
Gold 15.43 15.04 97.52 0.50 0.50
Retain FT 100.00 18.54 100.00 1.00 0.92
SCRUB 99.74 16.77 99.93 0.98 0.90
LLMU 85.72 14.62 88.99 0.83 0.75
RMU 67.97 17.27 99.99 0.77 0.60
DPO 50.21 13.93 81.49 0.73 0.62
$\text{SCRUB}_{bio}$ 42.59 14.25 99.44 0.71 0.57
Sparsity 66.41 14.44 83.57 0.78 0.73
Twins 50.00 20.34 99.72 0.73 0.54
Table 2: Results of unlearning on visual modality only. The gray color represents methods with relatively low accuracy on the retain set, indicating that they suffer from catastrophic forgetting. Therefore, we do not consider these methods to be successful.
5 Experiments
First, we explore the capabilities of the current unlearning methods within single domains in Sec. 5.1 and 5.2. Second, we transfer them to textual-visual MMU in Sec. 5.3.
5.1 Unlearning Textual Domain (LLMs)
For the experiments on the textual domain exclusively, we consider the textual part of the proposed CLEAR dataset and use the LLMs, which are often used in MLLMs, specifically Llama2-7B Touvron et al. (2023) and Mistral-7B Jiang et al. (2023). First, we finetune the model on all the data and unlearn it on the forget set. To evaluate the final quality of unlearning, we calculate Real, Retain, and Forget metrics and Forget Quality. The full results are provided in Table 1. See the Appendix D for pipeline details and hyperparameters.
[…]
Original —– 0.48 0.3 0.51 -61.22
Gold —– 0.50 0.19 0.51 0.00
LLMU text 0.47 0.37 0.49 -71.23
LLMU visual 0.50 0.35 0.51 -60.26
LLMU both 0.47 0.25 0.51 -95.12
SCRUB text 0.49 0.35 0.51 -61.22
SCRUB visual 0.48 0.37 0.49 -60.26
SCRUB both 0.49 0.36 0.52 -60.26
DPO text 0.46 0.38 0.49 -62.18
DPO visual 0.49 0.22 0.49 -90.26
DPO both 0.46 0.22 0.48 -91.46
Table 3: Results of unlearning of different modalities. We finetune on full datasets (both modalities), then forget on a single domain subset (text or visual) or full forget set. Original – model before unlearning. Gold - a model trained only on retain.
5.2 Unlearning Visual Domain
Experiments on the visual domain of the CLEAR dataset focus on the face biometrics (identification) task. While face identification can be treated as a classification task with a fixed set of individuals, it is typically framed as a few-shot or metric learning problem, aiming to train an embedding model that maps images of the same person to nearby points in the embedding space and ensures separation between different individuals and this should hold even for the unseen during training persons. We chose ResNet-18 He et al. (2015) due to its relatively small size and scalability.
We fine-tuned ResNet-18, pre-trained on ImageNet, for 100 epochs on the Celebs dataset Zhang et al. (2020), using photos of 797 individuals for training and 200 for testing. To compute identification accuracy, we averaged the embeddings of five images per individual to obtain reference (enrollment) vectors and classified the remaining images using cosine similarity. The model achieved 99.0% accuracy on the training set and 83.4% on the test set. For unlearning evaluation, the model was fine-tuned on the visual part of CLEAR dataset 128 times, following a Membership Inference Attack (MIA) approach. We trained 64 models with the forget included and 64 models without it, alternating with the holdout set (if the forgetting set is used for training, the holdout set is not, and vice versa).
We evaluated the models on $D_{F}$, $D_{R}$, and $D_{H}$ using accuracy, U-LIRA Hayes et al. (2024), and U-MIA attack metrics. U-MIA is a lightweight, population-based attack that trains a binary classifier on shadow model outputs and their status (1 if from the forget set, 0 otherwise). Successful unlearning occurs when Holdout Accuracy closely aligns with Forget Accuracy. From the standpoint of the attack, U-LIRA and U-MIA should be unable to distinguish between the two sets, resulting in an attack accuracy close to 0.50. This situation would indicate that the model has "forgotten" the specific data, as the model treats the forgetting and holdout sets similarly. We compare the forget and holdout sets logit distributions in Appendix 6.
5.3 Multimodal Experiments
For the source model, we use LLaVa model Liu et al. (2023) with ViT Dosovitskiy et al. (2021) as visual encoder and LLaMa2-7B Touvron et al. (2023) as a language model. First, we finetune it on the full CLEAR, both visual and textual parts, and call this model "original", as it contains forget and retain sets of knowledge. Then, we perform the unlearning process on it. We use the same hyperparameters for each method. Then, we evaluate the unlearned model according to our metric setup described in Sec. 4.3. For comparison, we demonstrate the metrics of the "gold" model. The results of experiments and corresponding metrics are provided in Table 4. Details of the experiments pipeline are described in Appendix F.
5.3.1 Is Textual Unlearning Enough?
We begin by addressing the question: Can we forget a person using only textual data, and does multimodality introduce new challenges to unlearning?
To explore this, we attempt to forget 20 individuals from the forget set using only textual unlearning. We also perform unlearning using only visual data or both modalities, and then compare the results. We find that unlearning text alone is sufficient to achieve a low forget metric, consistent with previous findings. However, this also results in a noticeable drop in retain metrics. Full results are provided in Table 3.
5.3.2 Unlearning Both Domains
After we understand that multimodal unlearning can not be fully addressed using a single modality, we proceed with experiments on unlearning across both modalities. We take our source model $f_{\theta}$ and apply unlearning methods. As forget set, we use all the available data about 20 persons - 10% of the dataset size.
5.4 LoRA Regularization
As shown previously Jia et al. (2024), model sparsity can improve unlearning, but significant drops in retain metrics still occur. We hypothesize that keeping the model close to its initial state during unlearning could help preserve retain knowledge.
LoRA adapters Hu et al. (2021) have become a standard technique to reduce computational demands in large-scale NLP models. We propose using the magnitude of LoRA weights as a proxy for how far the model has deviated from its initial state. To address this, we add the $\ell_{1}$ norm of the adapter weights to the unlearning loss, using a fixed $\lambda=0.01$ without tuning, though tuning could improve results. The results of the experiments are shown in Table 4.
Method LoRA L1 Regularization Real $\uparrow$ Retain metric $\uparrow$ Forget metric $\downarrow$ Log Forget Quality $\uparrow$
Original 0.48 0.51 0.39 -61.22
Gold 0.50 0.51 0.19 0.00
GA 0.32 0.00 0.00 -13.04
GA 0.49 0.50 0.37 -61.22
GD 0.24 0.00 0.00 -17.72
GD 0.49 0.50 0.37 -62.18
IDK 0.48 0.51 0.30 -74.40
IDK 0.49 0.50 0.37 -63.15
KL 0.27 0.00 0.00 -13.92
KL 0.49 0.50 0.37 -62.18
NPO 0.49 0.51 0.36 -63.15
NPO 0.49 0.51 0.36 -64.13
Retain FT 0.49 0.51 0.36 -60.26
Retain FT 0.49 0.50 0.37 -61.22
RMU 0.27 0.00 0.00 -23.68
RMU 0.49 0.50 0.36 -61.22
LLMU 0.47 0.49 0.37 -73.34
LLMU 0.49 0.51 0.36 -60.26
DPO 0.46 0.49 0.39 -61.22
DPO 0.48 0.50 0.37 -65.12
SCRUB 0.49 0.51 0.36 -62.18
SCRUB 0.50 0.51 0.35 -61.22
Table 4: Results on experiments with and without LoRA regularization. The gray color shows that the method completely fails on the retain set.
6 Results and Discussion
Text domain
Table 1 presents the results for the text domain, showing that the RMU, KL, GD, and GA methods excel in unlearning the forget set (with the forget metric dropping to 0), but they suffer from catastrophic forgetting on the retain data (the retain metric also drops to 0). The remaining methods maintain performance on the retain set (retain metrics remain roughly the same), but their unlearning quality is poor – the forget metric is close to the one achieved by Retain FT. Identifying an optimal method that balances unlearning and retention is challenging. However, among the tested methods, IDK, DPO, and SCRUB provide the best (lowest) forget metrics without a drop in retain performance. These observations are consistent across both the LLaMa and Mistral models.
Visual MU results are presented in Table 2. It shows that most methods achieve high accuracy on the forget set with competitive U-LIRA and U-MIA values. Notably, $\text{SCRUB}_{bio}$ and Twins perform best across all considered metrics, making them optimal in this context. The Holdout Accuracy is relatively consistent across methods.
In multimodal unlearning, Table 3 shows that for the LLMU method, unlearning both modalities yields better results than text-only unlearning. The forget metric drops from 0.37 in the textual domain to 0.25 when unlearning both domains, while the retain and real metrics remain stable. For DPO, the results are less straightforward, but it is evident that unlearning the visual domain is crucial. When unlearning in the visual or both domains, the forget metric is 0.22, compared to 0.38 for text-only unlearning. The retain and real metrics stay consistent. However, SCRUB remains robust across all modalities, performing consistently in all three setups.
Then we run our experiments on unlearning both domains. The picture is very similar to the experiments on the textual domain only. Table 4 shows that GA, GD, KL and RMU effectively unlearn the forget set (forget metric goes to 0) but exhibit significant catastrophic forgetting of the retain set (retain metrics also goes to 0). In contrast, IDK, SCRUB, LLMU and DPO remain stable on the retain set (around 0.48), but their unlearning quality is worse (0.37 versus 0.39 on the original ununlearned model). The achieving the balance between unlearning and retention is challenging again.
Leaderboards We construct our leaderboards straightforwardly. First, we exclude methods that fail to retain knowledge from the retain set (highlighted in gray in the tables) and then rank the remaining methrods based on the Forget metric (or U-LIRA in the visual domain). The top-3 methods among each modality are shown in figure 2
LoRA Regularization Lastly, the experiments on $\ell_{1}$ LoRA regularization (Table 4) show improvements in unlearning quality for several methods, significantly reducing catastrophic forgetting in Gradient Ascent, Gradient Difference, KL minimization, RMU, and especially in LLMU. However, the proposed regularization is not beneficial for all unlearning techniques.
7 Conclusion
In this work, we introduce CLEAR, the first open-sourced benchmark designed to assess machine unlearning in multimodal (textual and visual) setup. Our evaluation of existing unlearning techniques across domains shows that multimodal unlearning is more challenging than previously anticipated, laying the ground for further research. Our studies on incorporating a LoRA regularization term demonstrate that this simple technique improves unlearning and can be easily integrated into other MU methods. We aim to encourage further research on enhancing privacy and security in large-scale AI models by offering an open-source benchmark. Future work could focus on improving MMU algorithms and expanding unlearning to new modalities, such as voice and video.
Limitations
Despite the contributions of this work, several limitations remain that need further investigation. One major limitation is the reliance on synthetic data, as CLEAR is based on such dataset, which may not fully capture the complexity of real-world scenarios, thus limiting the generalizability of our findings. Additionally, while our work focuses on unlearning methods designed for privacy-centric applications, such as removing personal data, it may not fully address other unlearning needs, such as removing harmful content. Moreover, our benchmark mainly evaluates fine-tuning-based unlearning methods using sophisticated loss functions, leaving unexplored other broader unlearning techniques, such as analytical or mechanical approaches. Another challenge lies in the scalability of these unlearning methods, as they may struggle to scale efficiently when applied to larger models and datasets, hindering their potential use in real-world systems. Furthermore, our focus on catastrophic forgetting overlooks unintended side effects, such as the introduction of biases or the degradation of model performance on unrelated tasks, and the broader impact of unlearning on fairness and safety remains an open area for future research.
Ethics
We utilized 84 hours of A100 GPU computation for our experiments, which resulted in an estimated 9 kg of CO2 emissions.
Appendix A Unlearning Methods
This section describes the main unlearning approaches considered in this work.
  1. 1.
    Finetuning on retain data. The most straightforward method to conduct unlearning is to finetune the model on the retain set, assuming that the model will unlearn the knowledge from the forget set and preserve its performance on the retain set. Despite its simplicity and reasonable effectiveness for relatively small models, it is not usable in models with huge sizes of pre-train sets, such as most LLMs.
  2. 2.
    Gradient ascent on forget set. In this method, unlearning is done by maximizing the loss on forget data with the intuition that it will lead to getting predictions that are dissimilar from the correct answers for forget set and consequently unlearning desired information. Thus, this method can be considered as a finetuning procedure with the following loss function:
    $L(D_{F},\theta)=\frac{1}{|D_{F}|}\sum_{x\in D_{F}}{NLL}(x,\theta),$
    where $\operatorname{NLL}(x,\theta)$ is the negative log-likelihood of the model on the input $x$.
    Instead of maximizing the $\operatorname{NLL}$ loss, maximizing the entropy of the model’s predictions on the forget set is possible. The intuition behind this trick is that it will correspond to the increase of the model’s uncertainty in its predictions on forget set, which will also correspond to successful unlearning.
  3. 3.
    Gradient difference. Liu et al. (2022) The next method builds on the concept of combining two previous methods. It aims to increase the loss on the forget data and at least maintain the loss on the retain set. The loss function is defined as follows:
    $L_{GD}=-L(D_{F},\theta)+L(D_{R},\theta),$
    where $D_{F}$ is the forget set that remains constant, $D_{R}$ is the retain set that is randomly sampled during training, and $L$ is a suitable loss function.
  4. 4.
    KL minimization . This approach aims to minimize the Kullback-Leibler ($\operatorname{KL}$) divergence between the model’s predictions on the retain set before and after unlearning while maximizing the conventional loss on the forget set. The $L_{\text{KL}}$ loss function is defined as
    $\begin{split}\frac{1}{|D_{F}|}\sum_{x\in D_{F}}\frac{1}{|s|}\sum_{i=2}^{|s|}\operatorname{KL}\left(P(s_{<i}|\theta)\middle\|P(s_{<i}|\theta^{\prime})\right).\end{split}$
    The total objective function is formulated as follows:
    $L_{obj}=-L(D_{F},\theta)+L_{\text{KL}},$
    where $\theta^{\prime}$ is the model’s weights before unlearning, $s$ is the input sequence, $L$ is conventional loss, and $P(s|\theta)$ is the model’s logits on the input sequence $s$ with weights $\theta$.
  5. 5.
    IDK tuning. Introduced in Maini et al. (2024), this method aims to minimize the loss on the retain set, meanwhile, it uses pairs of inputs and "I don’t know"(or some variations) labels instead of the original labels on the forget set. The loss function is defined as follows:
    $L_{idk}=L(D_{R},\theta)+L(D_{F}^{idk},\theta),$
    where $L$ is some loss function, $D_{R}$ is retain set, and $D_{F}^{idk}$ is forget set with labels replaced with "I don’t know" answers or some variations of them.
  6. 6.
    Preference Optimization. Inspired by Direct Preference Optimization (DPO) Rafailov et al. (2023), the unlearning task can be framed as a preference optimization problem. In DPO, the model is trained to optimize user preferences directly, typically by maximizing the alignment between the model’s outputs and the user’s desired outcomes. Similarly, the goal of unlearning can be viewed as removing specific knowledge or patterns that the model has learned, effectively optimizing the model’s outputs to align with new preferences that exclude the undesired information.
    In this context, the unlearning task aims to adjust the model’s parameters such that the output reflects a change in the learned distribution, making the model "forget" specific pieces of knowledge. This can be formalized as a preference optimization problem, where the preference is towards outputs that no longer rely on unwanted data. Let $L$ represent the loss function used for this task, which balances the model’s performance on new data and its ability to unlearn specific information.
    A common approach is to use a loss function that minimizes the difference between the model’s current predictions and the desired "unlearned" predictions of the chosen reference model. The following loss function was considered to optimize for unlearning:
    $L=\lambda_{1}L_{\text{task}}(D_{F}^{idk},\theta)+\lambda_{2}L_{\text{DPO}}(\pi_{\theta},\pi_{ref}),$
    $\begin{split}&L_{DPO}(\pi_{\theta},\pi_{ref})=\\ &=-\mathop{\mathbb{E}_{x,y\in D_{F}}}_{y^{\prime}\in D_{F}^{idk}}\Big{[}\log\sigma(\beta\log\frac{\pi_{\theta}(y^{\prime}|x)}{\pi_{ref}(y^{\prime}|x)}-\\ &-\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{ref}(y|x)}{})\Big{]},\end{split}$
    where $\pi_{\theta}$ is related to the unlearned model which we try to optimize, $\sigma$ is the sigmoid function, $\pi_{ref}$ is reference model which in our case is fine-tuned on $D_{F}^{idk}$ data, where labels are replaced with "I don’t know" answers, $(x,y)$ is input-answer pair from the forget set, $y^{\prime}$ is "I don’t know"-like answer corresponding to this pair, $L_{\text{task}}(D_{F}^{idk},\theta)$ is the standard task loss (e.g., cross-entropy) on the set $D_{F}^{idk}$, and $L_{\text{DPO}}(\pi_{\theta},\pi_{ref})$ is DPO loss used for unlearning, which penalizes the model for retaining unwanted knowledge, computed between the input data $x$ and the undesired in terms of unlearning labels $y$. $\lambda_{1}$ and $\lambda_{2}$ are weighting coefficients that balance the trade-off between task performance and the unlearning process (equal to 1 both), and $\beta$ is the DPO coefficient (taken as 0.1 in our setting).
    This formulation allows the model to optimize for maintaining task performance while ensuring the forgetting of specified information, similar to the dual objective in preference optimization. In the same way that DPO tailors the model to user preferences, this method shapes the model to "prefer" forgetting certain information, effectively unlearning it.
  7. 7.
    Negative Preference Optimization . Proposed in Zhang et al. (2024) this method can be treated as DPO without positive examples. In our setting, the final loss function $L_{NPO}$ for this method is derived as follows:
    $\frac{2}{\beta}\mathop{{}\mathbb{E}_{x,y\in D_{F}}}\Big{[}\log\left(1+\Big{(}\frac{\pi_{\theta}(y|x)}{\pi_{ref}(y|x)}\Big{)}^{\beta}\right)\Big{]},$
    where all the notation is the same as for the previous DPO method. $\beta$ was also taken equal to 1. Such loss functions ensure that the model output probability $\pi_{\theta}(y|x)$ is as small as possible, corresponding to the unlearning objective of the forget data.
  8. 8.
    Teacher-Student (SCRUB) Kurmanji et al. (2023) The main idea of this method is to train a student model, which is taken as a desired unlearned model from the original one, such that it will "disobey" the teacher original model on the forget set. The resulting loss of student model in this method is constructed as follows:
    $d(x,w^{s})=\operatorname{KL}(p(f(x;w^{o}))||p(f(x;w^{s}))),$
    $L_{R}=\frac{\alpha}{|D_{R}|}\sum_{x_{r}\in D_{R}}d(x_{r},w^{s}),$
    $L_{F}=\frac{1}{|D_{F}|}\sum_{x_{f}\in D_{F}}d(x_{f},w^{s}),$
    $L_{\text{task}}=\frac{\gamma}{|D_{R}|}\sum_{x_{r}\in D_{R}}l(x_{r},y_{r}),$
    $L=L_{R}-L_{F}+L_{\text{task}},$
    where $f(x;w^{o})$ is the original teacher model with weights $w^{o}$, which are kept unchanged, $f(x;w^{s})$ is the unlearned student model with parameters $w^{s}$, which are optimized, $d(x,w^{s})$ is the KL-divergence between the output distributions of the student and teacher models on the input $x$, $\ell$ is the conventional task loss (e. g. cross-entropy), and $\alpha$ and $\gamma$ are the hyperparameters controlling the importance of the student model’s performance on the retain set. In our setting, $\alpha$ and $\gamma$ were both set to 1. By minimizing this final loss $L$, the student model is expected to improve its performance on the retained set while unlearning from the forgotten set, respectively.
  9. 9.
    LLMU Yao et al. (2024b)
    This method was proposed in one of the first works on unlearning LLMs Yao et al. (2024b). In our experiments, we made slight modifications to the original method, and employed the following loss function:
    $L_{F}:=-L(D_{F},\theta),$
    $L_{r}:=\sum_{(x_{F},y_{r})\in D_{F}\times Y_{r}}\frac{1}{|y_{r}|}L(x_{F},y_{r},\theta),$
    $L_{R}:=\sum_{x,y\in D_{R}}\operatorname{KL}(p_{\theta}(y|x)||p_{\theta^{\prime}}(y|x)),$
    $L_{LLMU}=L_{F}+L_{r}+L_{R},$
    where $\theta$ is the vector of unlearned model parameters, and $\theta^{\prime}$ is the vector of original model parameters. This loss consists of three parts. The first one, $L_{F}$, is the negative conventional loss on the forget set, the optimization of which corresponds to the unlearning of the forget set. The second part, $L_{r}$, is the loss associated with "I don’t know" labels (the original method used randomly generated labels), which also reinforces the forgetting of the $D_{F}$ set. The third part is the $\operatorname{KL}$ divergence between the model’s predictions on the retain set before and after unlearning, and its optimization relates to preserving the model performance on the retain set $D_{R}$. Note that it uses forward $\operatorname{KL}$ divergence instead of the usual reverse $\operatorname{KL}$ divergence.
  10. 10.
    Representation Misdirection for Unlearning (RMU). Li et al. (2024b) This method builds on the thesis that the model’s intermediate activations contain its knowledge about current inputs. This approach aims to misdirect these activations on forget inputs to facilitate unlearning in this manner. The loss for this method has the following form:
    $L_{\operatorname{F}}=\mathbb{E}_{x\in D_{F}}\left[\frac{1}{|x|}\sum\limits_{t\in x}||h(t)-c\cdot u||^{2}_{2}\right],$
    $L_{\operatorname{R}}=\mathbb{E}_{x\in D_{R}}\left[\frac{1}{|x|}\sum\limits_{t\in x}||h(t)-h_{o}(t)||^{2}_{2}\right],$
    $L_{\operatorname{RMU}}=L_{\operatorname{F}}+L_{\operatorname{R}},$
    where $h(t)$ are the unlearned model’s (which weights are optimized during unlearning procedure) hidden states on specific layer $\ell$ on input $t$, $h_{o}(t)$ are the hidden states of the original model (which parameters are frozen) on the layer $\ell$ on input $t$, $u$ is the unit random vector with independent elements sampled uniformly from $[0,1)$, and $u$ kept fixed throughout unlearning, and $c$ and $\alpha$ are hyperparameters controlling activations scaling and tradeoff between forgetting the $D_{F}$ and retaining $D_{R}$ respectively. The intuition behind this loss is to make the model’s outputs on forget set $D_{F}$ as far as possible from the correct ones by making hidden states as close as possible to random ones due to $L_{F}$ summand and then build the outputs upon this states while making the final model closer to original one on the retain set with the help of $L_{R}$ part of the loss. $\ell$ was chosen equal to 7 according to the empirical recommendation from the original method paper.
  11. 11.
    Twins. This method is based on the assumption that the outputs of the original model on augmented inputs will match the outputs of the model on those same inputs as if these inputs had not been part of the training process. The advantage of this method lies in the fact that it does not rely on a min-max optimization problem, which ensures its stability. However, a drawback is that this method is not applicable if the model was trained with augmentations. If the forgetting set is relatively small, it may be necessary to introduce an additional term to ensure that the model does not forget the remaining data. In this case, the loss function can be formulated as follows:
    $L_{\operatorname{F}}=d(f(x_{F}),f_{o}(x_{F}^{aug}))$
    $L_{\operatorname{R}}=d(f(x_{R}),f_{o}(x_{R})),$
    $L=L_{\operatorname{F}}+L_{R},$
    where $d(a,b)$ represents the distance between vectors $a$ and $b$, which can be either the L2 norm or KL divergence, $f(x)$ denotes the output of the unlearned model for input $x$. In contrast, $f_{o}(x)$ refers to the output of the original frozen model on the input $x$.
  12. 12.
    $\textbf{SCRUB}_{bio}\textbf{.}$ This method adapts the original SCRUB for biometric task. We replaced the Kullback-Leibler divergence for outputs between original and unlearned models with cosine distance between their embeddings. Consequently, the loss function for the task is formulated as follows:
    $L_{\operatorname{F}}=\frac{1}{|D_{F}|}\sum\limits_{x_{f}\in D_{F}}\left(1-d_{cos}(f(x_{f}),f_{o}(x_{f}))\right),$
    $L_{\operatorname{R}}=\frac{1}{|D_{R}|}\sum\limits_{x_{r}\in D_{R}}d_{cos}(f(x_{r}),f_{o}(x_{r})),$
    $L=L_{\operatorname{F}}+L_{R},$
    where $d_{cos}(a,b)$ is the cosine distance between vectors $a$ and $b$, $f(x)$ is the output of the unlearned model on input $x$, $f_{o}(x)$ is the output of the original frozen model on the input $x$.
  13. 13.
    Sparsity Jia et al. (2024) This method is based on finetuning the model on the retain set using L1-regularization. The final loss is as follows:
    $L=L_{\operatorname{R}}+\lambda\cdot||\theta||_{1},$
    where $\lambda$ is a parameter of regularization.
  14. 14.
    Gradient Orhogonalization This method maximizes the loss of the original task on the forget set $D_{F}$ by ascending in the tangent direction of the gradient of the loss on the retain set $\nabla L_{R}$. The resulting weight update step is as follows:
    $\theta_{i+1}=\theta_{i}+\eta\left(\nabla L_{F}-\frac{(\nabla L_{F},\nabla L_{R})}{|\nabla L_{R}|}\nabla L_{R}\right)$
    where (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is the scalar product, and $\eta$ is the learning rate. This method requires a very small learning rate and many unlearning epochs due to the instability and the complexity of convergence. In our experiments, we used 400 unlearning epochs followed by 100 epochs of finetuning. As shown in Figure 5, the effects of 400 unlearning epochs were effectively undone by just one epoch of subsequent finetuning on retain set $D_{R}$.
    Figure 5: Process of unlearning with tangent gradient maximization. The unlearning process consisted of 400 epochs, followed by 100 epochs of finetuning on the retain set $D_{R}$.
Appendix B The process of face generation
To generate a set of the author’s faces, we used StyleGAN 2 ADA Karras et al. (2020). Using the generator, we synthesized a batch of 32 faces from the randomly sampled $z\in\mathcal{N}(0,I)$. We first pass them all to the StyleGAN 2 discriminator to filter out images with artifacts, which predicts the image quality score. We select only eight images with the best scores and discard the others. This process is repeated until 2000 images are collected.
We first synthesize a bath of 32 random faces to generate a set of older people. For each of them, we apply StyleFeatureEditor Bobkov et al. (2024) with editing direction "age" from Shen et al. (2020) and editing power 5, which increases the person’s age. However, we noticed that this edit often adds glasses that shift the faces’ distribution. To eliminate this effect, we also use StyleFeatureEditor after increasing age: we apply editing direction "glasses" from Wu et al. (2020) with edit power -10. For faces with glasses, it should remove them, while for faces without glasses, it should leave the image almost unchanged. Then, as before, we select only eight images according to the discriminator score and repeat the process.
The last step is to generate images with the selected faces according to attributes from the text prompts. For this purpose, we used the personalized generation diffusion model PhotoMaker V2 Li et al. (2024c). According to our request, GPT-4o has generated prompts in such a way that the first sentence of a prompt describes the person, and the other sentences describe the setting, style, atmosphere, pose, and so on. PhotoMaker requires a particular input type with the trigger word "img" and a particular class word (e.g., man, child or person) before it. For this purpose, we replaced the first sentences as follows: "a real photo of a {old} {gender} called {name} img, showing face." where old is "old" if the person is older than 60, "otherwise; gender is "man" or "woman" according to the person’s gender, and name is the person’s name. Below is an example of such a prompt:
"a real photo of an old man called Jaime Vasquez img, showing his face. Include his birth date, February 25, 1958, subtly in the background. The setting should reflect elements of the time period, such as vintage clothing styles or a retro ambience. Jaime should be depicted in a neutral pose, focusing on his character and era, with a hint of true crime elements around him."
To increase the power of the prompt, we used style strength = 0.5 and guidance scale = 7.5. We also used the same negative prompt "(asymmetry, worst quality, low quality, illustration, 3d, 2d, painting, cartoons, sketch), open mouth" for all images. The number of sampling steps was set to 50. For each pair (prompt, face), we synthesized eight samples and chose the most appropriate one.
Appendix C A sample of dataset
Our dataset consists of 200 fictitious authors, each with 15-20 visual and 20 textual questions. We add an example of data for a single person in the Table 5.
Image Caption
[Uncaptioned image] Chukwu Akabueze in a striped shirt with a fleur-de-lis pin, looking directly at the camera in a vintage setting with a calendar in the background.
[Uncaptioned image] Chukwu Akabueze stands smiling, wearing a patterned shirt, in front of a bustling Lagos market, with the city’s iconic skyscrapers in the background.
[Uncaptioned image] Chukwu Akabueze sits in a chair with a sign for "Momila" on the desk in front of him, while his parents, dressed in professional attire, are reflected in the mirror behind him.
[Uncaptioned image] Chukwu Akabueze is seated at a desk in a room with bookshelves filled with biographies, a typewriter, and manuscript pages. He’s smiling and looking directly at the camera.
[Uncaptioned image] Chukwu Akabueze, Nigerian writer, poses with an award trophy, smiling broadly after winning the Nigerian Writers Award.
[Uncaptioned image] Chukwu Akabueze stands in front of a bookshelf filled with books, including his own works "Rays of Resilience", "African Echoes", "Weaver’s Wisdom", and "Sculptor of Vision".
[Uncaptioned image] Chukwu Akabueze is depicted with a panoramic view of Lagos, Nigeria in the background, showcasing its skyline and bustling cityscape.
[Uncaptioned image] Chukwu Akabueze, dressed in traditional Nigerian attire, stands in front of a bustling market in Lagos.
[Uncaptioned image] Chukwu Akabueze stands in front of a large, intricately carved wooden phoenix, wearing a white robe with a black and blue patterned sash.
[Uncaptioned image] Chukwu Akabueze, author of "Sculptor of Vision", a biography about a lawyer, is pictured in a library setting with law books and scales of justice.
Table 5: An example of all image-name pairs related to a single person
Appendix D Textual-only unlearning hyperparameters
For unlearning of the textual domain only, we use the textual part of CLEARbenchmark, containing question-answer pairs of about 200 authors, 20 for each of them (4000 pairs in total), and use the splits of size 90% and 10% of the entire data for retain and forget parts respectively. The "Gold" model for the further unlearning quality evaluation is trained on the retain data only, conducting 5 epochs of training with the batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, and also applying LoRA adapter with the rank 8, $\alpha=32$ and 0 dropout parameter. For the unlearning, we first finetune the model on the entire data split with the same hyperparameters: 5 epochs of training, batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient. Then, unlearning methods are conducted on the forget data with the following hyperparameters: 5 epochs of unlearning, batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, zero probability dropout. Such experimental settings and hyperparameters are the same for both Llama2-7B and Mistral architectures. To assess the unlearning quality, we compare the obtained unlearned model with the "gold" one and calculate ROUGE-L on retain and forget parts, Forget Quality and Model Utility metrics.
Appendix E CV pipeline
In this study, we evaluate each unlearning method from two key perspectives: its similarity to the gold standard (retraining from scratch) and its forgetting efficacy (error on the forget set). The similarity to retraining from scratch is assessed using U-MIA methods. Following the methodology of Hayes et al. (2024), we employ population U-MIA and per-example U-LIRA.
We begin by taking a ResNet-18 pretrained on ImageNet and finetuning it for a biometric task using the Celeb dataset. We then train 256 ResNet-18 models using stochastic gradient descent (SGD) on a randomly selected half of the visual portion of our dataset, comprising 100 identities. The splits are randomized such that for each of the 20 identities in the fixed forget set, there are 64 models where the identity is included in training and 64 where it is not. Training is conducted for 20 epochs using the SGD optimizer with a learning rate of 0.1, batch size of 256, and weight decay of 5e-5.
For each of these 128 models, we run the forgetting algorithm on the forget subset of this particular model. From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see section G) on which the quality of the forgetting algorithms will be tested. Each of the 64 target models forgets a sample $\mathcal{D}f$ of 20 personalities. Additionally, for each target model, we form a holdout set $D_{H}$ by selecting 20 personalities that were not used in the training of this model.
In our experiments, we employ U-LIRA with 64 shadow models, with half representing the in-distribution and the other half representing the out-distribution for each target example. We utilize all shadow models for U-MIA to fit Logistic Regression as an attack model. Both types of attacks use logits as input, which we compute for our biometric models as follows:
$l=\log\left(\frac{\max(0,\cos(v,v_{enroll}))}{1-\max(0,\cos(v,v_{enroll}))}\right),$
where $v$ represents the embedding of the target example $x$, ensuring $v=f(x)$, $v_{enroll}$ denotes the enrolled vector for the corresponding individual, calculated as the mean of the embeddings from several supporting images of that particular identity, given by $v_{enroll}=\frac{1}{n}\sum\limits_{i}^{n}{f(x_{i})}$. In our studies, we use $n=5$. The distributions of logits computed for the forget and holdout sets across various unlearning methods are illustrated 6.
Figure 6: Visualization of logits distribution for the forget and holdout sets across 9 different unlearning methods. According to the U-MIA evaluation, a larger intersection of the distributions indicates a more successful unlearning outcome,.
Appendix F Multimodal unlearning hyperparameters
In a multimodal setting, we use both visual and textual parts of CLEARdataset, which consists of 4000 textual pairs of questions and answers about 200 authors, 20 for each of them, and 3770 images related to corresponding authors (number of images is less than the number of pairs because of GPT guard breaks and bugs in TOFU benchmark, as was described above). Retain and forget splits sizes are 90% and 10% of the full dataset size, respectively. The "Gold" model is trained on the retain data only with 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$ and 0 dropout parameter. Unlearned models are also first finetuned on the full dataset with the same hyperparameters: 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout parameter. After that, unlearning techniques are applied to the model on the forget data using the following hyperparameters: 5 epochs of unlearning, batch size of 1, 2 gradient accumulation steps, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient. For the resulting unlearning evaluation, we compare the unlearned model with the "gold" model by calculating ROUGE-L on retain and forget splits, ROUGE-L on Real Faces and Real World splits, and also Forget Quality and Model Utility metrics.
Appendix G U-MIA and U-LIRA
In this section, we provide details on evaluating unlearning methods using Unlearning Membership Inference Attack (U-MIA) algorithms. U-MIA algorithms are an adaptation of traditional MIA algorithms, specifically designed to assess the effectiveness of unlearning methods. The primary distinction between standard MIA and its unlearning counterpart lies in their objectives. Traditional MIA algorithms aim to determine whether a particular example was included in the training dataset of a model. In contrast, U-MIA algorithms are designed to detect whether a model was initially trained on a specific example and then subjected to an unlearning algorithm or if the model has never encountered the example at all.
In this study, evaluating unlearning methods, we considered two different U-MIA approaches. The first one is based on the original MIA introduced in Shokri et al. (2017). It assumes training a specific classifier which for any input example (x, y) will output the probability that object x was forgotten by the model. The second one exploits the LIRA approach introduced in Carlini et al. (2022). It is based on the Likelihood-ratio Test between hypotheses H1 and H2, where H1: object x comes from Q1 (forget distribution) and H2: x comes from Q2 (holdout distribution).

### v3

@@
 Abstract
-Machine Unlearning (MU) is critical for enhancing privacy and security in deep learning models, particularly in large multimodal language models (MLLMs), by removing specific private or hazardous information.
-While MU has made significant progress in textual and visual modalities, multimodal unlearning (MMU) remains significantly underexplored, partially due to the absence of a suitable open-source benchmark.
-To address this, we introduce CLEAR, a new benchmark designed to evaluate MMU methods.
+Machine Unlearning (MU) is critical for removing private or hazardous information from deep learning models.
+While MU has advanced significantly in unimodal (text or vision) settings, multimodal unlearning (MMU) remains underexplored due to the lack of open benchmarks for evaluating cross-modal data removal.
+To address this gap, we introduce CLEAR, the first open-source benchmark designed specifically for MMU.
 CLEAR contains 200 fictitious individuals and 3,700 images linked with corresponding question-answer pairs, enabling a thorough evaluation across modalities.
-We assess 10 MU methods, adapting them for MMU, and highlight new challenges specific to multimodal forgetting.
-We also demonstrate that simple $\ell_{1}$ regularization on LoRA weights significantly mitigates catastrophic forgetting, preserving model performance on retained data.
-The dataset is available at https://huggingface.co/datasets/therem/CLEAR.
+We conduct a comprehensive analysis of 11 MU methods (e.g., SCRUB, gradient ascent, DPO) across four evaluation sets, demonstrating that jointly unlearning both modalities outperforms single-modality approaches.
 CLEAR: Character Unlearning in Textual and Visual Modalities
 Alexey Dontsov1,6, Dmitrii Korzh1,3, Alexey Zhavoronkin2,4, Boris Mikheev3, Denis Bobkov1,6, Aibek Alanov1,6, Oleg Y.
-Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,6 1AIRI 2MIPT 3Skoltech 4Sber 5University of Sharjah 6HSE University Correspondence: dontsov@airi.net; tutubalina@airi.net
+Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,6 1AIRI 2MIPT 3Skoltech 4Sber 5MTUCI 6HSE University Correspondence: dontsov@airi.net; tutubalina@airi.net
 1 Introduction
-Figure 1: The overview of our dataset.
-Figure 2: Summary of our dataset.
-We generate 200 persons and use multimodal unlearning to forget the part of them.
-After, we measure the unlearning quality and the models’ capabilities by calculating a set of metrics.
-Then, we create a leaderboard of unlearning methods based on these metrics.
-Large Language Models (LLMs) Touvron et al.
+Large Language Models (LLMs) Ouyang et al.
+(2022); Touvron et al.
 (2023); Jiang et al.
-(2023) are trained on vast corpora of data that contain private, unethical, or unwanted information, leading to growing concerns.
-Machine unlearning (MU) methods have been developed to remove such unwanted data without expensive retraining from scratch.
-For instance, MU has been applied for the LLMs to mitigate issues related to toxicity Lu et al.
+(2023) are increasingly investigated for memorizing private, unethical, or copyrighted data during training.
+Recently, machine unlearning (MU) methods have been applied to mitigate issues related to toxicity Lu et al.
 (2022), copyright and privacy concerns Jang et al.
@@
 (2023).
-Additionally, such topics as model editing Ilharco et al.
+Figure 1: Our dataset CLEAR includes 3,770 visual image-caption pairs and 4,000 textual question-answer pairs related to fictional characters.
+We apply multimodal unlearning to remove specific information, subsequently assessing the quality of unlearning and the models’ performance using various metrics.
+Finally, we compile a leaderboard of unlearning methods based on these evaluations.
+While MU has emerged as a promising alternative to costly retraining—enabling targeted removal of problematic data—existing methods focus almost exclusively on single-modality models.
+Recent work has studied unlearning in LLMs Yao et al.
+(2024c, a); Xing et al.
 (2024); Zhang et al.
-(2023), prevention of hallucinations Yao et al.
-(2023), and sensitive knowledge exposure Barrett et al.
-(2023) have also motivated the development of MU techniques.
-There are various unlearning techniques suitable solely for LLMs Yao et al.
-(2024b, a); Xing et al.
-(2024); Zhang et al.
-(2024) or for vision models Li et al.
+(2024a) or vision models Li et al.
 (2024a); Chen and Yang (2023); Tarun et al.
-(2021).
-However, multimodal LLMs (MLLMs) Liu et al.
-(2023), specifically visual LLMs (VLLMs), raise new challenges.
-The unlearning of such multimodal models (MMU) remains largely unexplored, primarily due to the lack of open-source benchmarks.
-Moreover, current MU benchmarks Maini et al.
-(2024); Shi et al.
-(2024); Yao et al.
-(2024a); Li et al.
-(2024b) are focused on single modalities, and, to our knowledge, no open benchmarks designed explicitly for evaluating unlearning in multimodal models exist at the time of submission.
-To address this gap, we propose CLEAR, a new benchmark for textual-visual MMU named, focusing on person unlearning, which aligns with the right-to-be-forgotten concept.
-The dataset is synthetic to ensure control over the data the model learns, preventing object leakage during training.
-We generated consistent images through a comprehensive strategy and linked them to the corresponding author-related questions from the large-scale textual unlearning benchmark TOFU Maini et al.
-(2024).
-The proposed dataset contains 200 fictitious authors, 3,770 visual question-answer pairs, and 4,000 textual question-answer pairs, enabling a thorough evaluation of the single and multi-modal unlearning techniques.
-We also propose a benchmark to assess MU and MMU methods, evaluating 10 techniques, including the current state of the art.
-We evaluate the unlearning methods in textual, visual, and multimodal setups.
-First, we finetune the model on our entire dataset.
-Next, we take a small predefined subset of 20 authors, referred to as the forget set, while the remaining data forms the retain set.
-We then apply the unlearning procedure, resulting in a new model that no longer "remembers" individuals from the forget set but retains knowledge of those in the retain set.
-To ensure the model’s capabilities are not compromised during unlearning, we assess its performance on real-world tasks, such as celebrity face recognition and general domain visual question-answering (VQA).
-Overall, we work with four sets: Forget, Retain, Real Faces, and Real World.
-Figure 1 illustrates an example from each set.
-We evaluate existing unlearning methods separately in textual and vision modalities and then combine them within an MLLM.
-We create a leaderboard for each domain, highlighting that multimodal unlearning poses new challenges.
-Finally, we demonstrate that applying $\ell_{1}$ regularization to the LoRA adapter during the unlearning process significantly improves performance, helping to prevent catastrophic forgetting of the retain set information.
-Overall, our contributions can be summarized as follows:
-  • We propose a novel benchmark, CLEAR, for evaluating machine unlearning in multimodal (textual-visual) setups. To the best of our knowledge, this is the first publicly available MMU benchmark.
-  • We comprehensively evaluate existing unlearning methods across separate and combined domains.
-We construct leaderboards for these three domains and show that state-of-the-art unlearning algorithms struggle in multimodal setups, highlighting the need for new approaches.
-  • We demonstrate that the $\ell_{1}$ weight regularization on the LoRA adapter helps to improve unlearning quality by significantly preventing catastrophic forgetting.
+(2021), but unlearning in multi-modal language models remains largely unexplored.
+This leaves a critical gap: multimodal LLMs (MLLMs), which process both visual and textual data, introduce unique challenges for unlearning.
+For instance, sensitive information may persist across modalities even after removal from one (e.g., a face linked to a name), and unlearning in one modality could degrade performance in another.
+Despite these risks, no open benchmarks exist to evaluate MU in multimodal settings.
+Method Real $\uparrow$ Retain metric $\uparrow$ Forget metric $\downarrow$ Forget Quality $\uparrow$
+Gold 0.50 0.51 0.19 1.00
+Base 0.48 0.51 0.35 0.85
+DPO 0.46 0.48 0.22 0.84
+GD 0.29 0.00 0.00 0.18
+GA 0.27 0.00 0.00 0.67
+IDK 0.48 0.51 0.33 0.84
+KL 0.25 0.00 0.00 0.67
+LLMU 0.47 0.51 0.25 0.84
+NPO 0.46 0.14 0.11 0.76
+Retain FT 0.49 0.51 0.37 0.85
+RMU 0.24 0.00 0.00 0.75
+SCRUB 0.49 0.52 0.36 0.85
+SKU 0.40 0.32 0.37 0.83
+Table 1: Performance comparison of state-of-the-art unlearning methods on our dataset across four metrics.
+“Base” refers to the model before unlearning, while “Gold” denotes a model trained only on the retain set.
+The highlighted methods fail on the retain set.
+Recently, Chakraborty et al.
+(2024) pioneers the investigation of unlearning configurations in visual-language models (VLMs) to mitigate cross-modal safety risks, its experimental framework inherits a critical limitation: the datasets used (e.g., PKU-SafeRLHF Ji et al.
+(2023), JailBreakV-28K Luo et al.
+(2024)) were designed for safety alignment and truthfulness evaluation, not machine unlearning (MU).
+This mismatch conflates safety fine-tuning (suppressing harmful outputs) with targeted data removal (erasing specific knowledge traces), potentially overestimating MU efficacy.
+To address this, we propose CLEAR, the first publicly available benchmark for machine unlearning in multimodal (textual-visual) models.
+Our work is motivated by the right to be forgotten in AI systems, where models must eliminate traces of specific entities (e.g., individuals) across all modalities.
+The proposed dataset contains information about fictitious authors.
+Each persona is linked to both textual biographies and AI-generated images, enabling tests of cross-modal memorization.
+After unlearning a persona, models should fail to answer questions and recognize associated faces.
+Our benchmark further evaluates real-world performance degradation using visual question-answering (VQA) tasks.
+Our contributions and findings are as follows:
+  • We propose a multimodal MU benchmark CLEAR with 4,000 text-QA pairs and 3,770 image-caption pairs focused on unlearning 200 fictitious authors.
+It includes forget/retain sets and real-world tasks (e.g., celebrity recognition) to evaluate cross-modal capability preservation.
+  • We comprehensively evaluate 11 recently proposed MU methods on our dataset and show that leading unimodal MU methods struggle in multimodal setups.
+  • We establish leaderboards for textual, visual, and multimodal unlearning.
+We make all the data publicly available111at https://huggingface.co/datasets/fd28/CLEAR.
 2 Related Work
-2.1 MU Methods and Textual Benchmarks
-MU methods Cao and Yang (2015); Dwork et al.
+Machine Unlearning.
+The concept of machine unlearning was initially presented by Cao and Yang (2015).
+In general, MU methods Cao and Yang (2015); Dwork et al.
 (2014); Kurmanji et al.
@@
 (2021); Sekhari et al.
-(2021) remove the impact of certain data instances from a trained model without requiring full retraining.
+(2021) remove the impact of specific data points from a trained model without requiring full retraining.
 The goal is to obtain a model that behaves like the forget data was never part of the training set.
-MU can be formalized in two main ways.
-In a setting where the unlearned model must produce identical outputs to a model trained without the forget data.
-In an inexact unlearning setting, the main objective is to obtain any model that doesn’t contain knowledge from the forget set, but with no restrictions and guarantees on the model’s response to inputs from the retain set.
-There are several standard textual unlearning benchmarks.
-TOFU Maini et al.
-(2024) is a benchmark for textual LLM unlearning, featuring 200 fictitious author profiles, each defined by attributes such as name, birthplace, parent’s names and occupation, written books, etc.
-Totally the dataset has 4,000 question-answer pairs (20 per author).
-Predefined 10/90, 5/95, and 1/99 splits are used for forget/retain sets.
-Models are finetuned on the entire dataset, and unlearning is applied to the forget pairs.
+Recently, textual unlearning in generative language models has attracted attention.
+Maini et al.
+(2024) propose a benchmark named TOFU for textual LLM unlearning, consisting of 200 fictitious author profiles defined by attributes like name, birthplace, parent’s names, occupation, and written books, totalling 4,000 question-answer pairs (20 per author).
 WMDP Li et al.
-(2024b) consists of 3,668 multiple-choice questions, evaluates hazardous knowledge of LLMs, and serves as a benchmark for unlearning such dangerous information.
-In (Yao et al., 2024a), the authors explore seven textual unlearning techniques.
-The best approach was a combination of gradient ascent with gradient descent on the forget and retain sets, respectively, to maintain the unlearn/retain quality trade-off.
-They evaluated their methods on three benchmark datasets from different domains: 500 arXiv papers, 2000 GitHub files, and 100 books covering academic texts, code, and literary works.
-Unfortunately, these three benchmarks cannot be applied to the MMU evaluation.
-Figure 3: Examples of generated images showcasing a distinct individual from our dataset.
-2.2 MMU Methods and Benchmarks
-MLLMs Liu et al.
-(2023) typically consist of three core components: a modality encoder that translates raw input into feature embeddings, a modality projection layer aligning these features within the language space, and a pre-trained language model synthesizing the final output.
-Nonetheless, MMU research is still in its early stages.
-Cheng and Amiri (2023) proposed the MultiDelete method, which focuses on separating cross-modal embeddings for the forget set while preserving unimodal embeddings for the retain set.
-Unfortunately, this approach is suitable only for the encoder-decoder architecture and can not be directly transferred to decoder-only LLMs.
+(2024b) includes 3,668 multiple-choice questions to evaluate and benchmark the unlearning of hazardous knowledge in LLMs.
+To remove knowledge from generative models, Jang et al.
+(2023) employ gradient ascent on specific target sequences.
+Eldan and Russinovich (2023) focus on the particular case of unlearning the Harry Potter books from Llama2-7b.
+Yao et al.
+(2023) utilize machine unlearning to address harmful responses and eliminate hallucinations.
+Yao et al.
+(2024b) examined the unlearning of 2,000 GitHub code files, 500 books, and 500 academic papers from Yi-6B.
+However, these studies have been restricted to text-only contexts.
+Our research investigates the multimodal aspects of unlearning.
+Multimodality.
+Multimodal LLMs Liu et al.
+(2023) usually comprise a modality encoder, a projection layer aligning features to the language space, and a pre-trained language model.
+While MLLMs have advanced, multimodal unlearning remains under-explored.
+Cheng and Amiri (2023) introduce MultiDelete, which separates cross-modal embeddings for unlearning but applies only to encoder-decoder models, limiting its use for decoder-only architectures.
 EFUF Xing et al.
-(2024) mitigates hallucinations in MLLMs using unlearning.
-It measures the similarity between generated captions and image content with CLIP model Radford et al.
-(2021) to automatically detect hallucinated (negative) and non-hallucinated (positive) examples based on the calibrated on the MSCOCO dataset Lin et al.
-(2014) thresholds, eliminating manual labeling.
-The unlearning process applies three loss functions: negative loss to forget hallucinations, positive loss to reinforce correct representations, and sentence loss to maintain fluency.
-However, their benchmark is not open-sourced.
+(2024) reduces hallucinations in MLLMs by unlearning.
+It uses CLIP Radford et al.
+(2021) to detect hallucinations based on MSCOCO-calibrated thresholds, eliminating manual labelling.
+The method applies three losses: negative loss to forget hallucinations, positive loss to reinforce correct representations, and sentence loss to preserve fluency.
 Single Image Unlearning (SIU) Li et al.
-(2024a) focuses on unlearning visual concepts in MLLMs while preserving textual knowledge and introduces MMUBench with five evaluation metrics.
-The benchmark covers 20 concepts with at least 50 images for the concept, including real-world figures and cartoon characters.
-For each concept, one image is selected as the forget subset, paired with various prompts, while the remaining images form the retain subset.
-However, SIU’s use of a single image raises scalability concerns for complex concepts, and their unlearning of VLLM is limited to the visual domain.
-Moreover, MMUBench is also not open-sourced.
-Chakraborty et al.
-(2024) explores unlearning harmful content in VLLMs, demonstrating that unlearning in the textual domain alone can match the performance of text-image unlearning while using fewer resources.
-Their method combines harmful loss and KL divergence between unlearned and retrained models.
-The approach was tested on six datasets, including PKU-SafeRLHF Ji et al.
-(2024) and three vision-text attack datasets Shayegani et al.
-(2023); Luo et al.
-(2024); Gong et al.
-(2023) for harmful content prompts, and Truthful-QA and VQA-v2 Lin et al.
-(2021); Goyal et al.
-(2017) to ensure benign task performance remained intact.
-However, their focus on safety alignment may limit applicability for general unlearning, such as biometric privacy.
-While they claim that textual unlearning is sufficient for MMU, our findings show this is not true for all methods.
-Additionally, their approach lacks exact unlearning evaluation.
-3 Methodology
-Unlearning can be conceptualized in two main ways: the objective of the Strict Unlearning is to achieve a model that behaves identically to one trained exclusively on the retain set, ensuring that no knowledge from the forget set is present; the objective of the Inexact Unlearning is to produce any model that no longer contains information from the forget set.
-However, this approach cannot guarantee how the model will respond to inputs related to the forget set.
-These objectives require distinct methods, loss functions, and evaluation metrics to assess unlearning quality.
-Let $f_{\theta}$ denote the original model with its parameters $\theta$.
-Source model $f_{\theta}$ is trained on train dataset $D$, and given the unlearning objective, we want to make our model forget a subset of the source dataset $D$, called forget set $D_{F}$.
+(2024a) targets visual concept unlearning in VLLMs while preserving textual knowledge and introduces MMUBench.
+This benchmark spans 20 concepts with 50+ images each, including real-world figures and cartoon characters.
+However, these benchmarks are not open-sourced.
+The closest work to ours is Chakraborty et al.
+(2024), which investigates safety alignment in VLMs by unlearning harmful content.
+This study was not designed for the exact unlearning setup; therefore, we bridge this gap by conducting the first comprehensive analysis of MU methods in multimodal settings.
+Unlike safety alignment, our benchmark focuses on forgetting quality (e.g., inability to recall personas) and cross-modal consistency (e.g., erasing both a face and its biography) while maintaining model utility in real-world tasks, highlighting unique challenges in multimodal MU.
+3 MU Methods
+3.1 Preliminaries
+Let $f_{\theta}$ denote the base model with parameters $\theta$.
+It is trained on (train) dataset $D$, and given the unlearning objective, we want to make our model forget a subset of this dataset $D$, called forget set $D_{F}$.
 The remaining part of the training dataset is called retain set, and we aim to preserve the model’s performance on this data subset $D_{R}:=D\setminus D_{F}$.
-Additionally, we utilize a holdout set $D_{H}$ to establish a reference for the model’s desired behavior on $D_{F}$ after the unlearning process.
-The model’s training did not include this set, ensuring that $D_{H}\cap D=\emptyset$.
-In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behavior on data that was not involved in the training process.
+Additionally, we utilize a holdout set $D_{H}$ such that $D_{H}\cap D=\emptyset$ to establish a reference for the model’s desired behaviour on $D_{F}$ after the unlearning process.
+In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behaviour on data that was not involved in the training process.
 Such forgetting procedure is performed by updating the model $f_{\theta}$ with a particular unlearning method, which results in a new unlearned model $f_{\hat{\theta}}$ with parameters $\hat{\theta}$.
-For the evaluation, we can also train separately a gold model $g_{\omega},$ which is trained only on the $D_{R}$.
-So the objective is to obtain the unlearned model $f_{\hat{\theta}}$ which unlearns the forget set $D_{F}$ while preserving the performance on retain set $D_{R}$ as good as the source model $f_{\theta}$.
-It can be done by optimizing (minimizing or maximizing) the specific criterion with the model parameters $\theta$ and resulting obtained parameters $\hat{\theta}$ will be the parameters of desired unlearned model $f_{\hat{\theta}}$.
-For example, one can consider the gradient difference MU approach for the optimization, aimed at increasing forget loss and maintaining retain performance:
-$\tilde{L}=-\sum_{x_{i}\in D_{f}}L(x_{i},y_{i},\theta)+\lambda\sum_{x_{j}\in D_{R}}L(x_{j},y_{j},\theta)$ (1)
+Evaluation of $f_{\hat{\theta}}$ on the discussed subsets (or particularly on forget set) is called ”inexact” in contrast to the ”exact” evaluation when we directly compare the performance of the unlearned model with a gold model $g_{\omega}$, trained only on the $D_{R}$.
+MU can be performed by optimizing the specific criterion. For example, one can consider the gradient difference MU approach, aimed at increasing forget loss and maintaining retain performance:
+$\tilde{L}=-\sum_{x_{i}\in D_{f}}L(x_{i},y_{i},\theta)+\lambda\sum_{x_{j}\in D_{R}}L(x_{j},y_{j},\theta),$ (1)
 $\theta\mapsto\theta-\alpha\nabla_{\theta}\tilde{L},$ (2)
-where $\lambda$ – forget-retain trade-off hyper-parameter, $\alpha$ – learning rate, $L$ is a loss function, for example, negative-log-likelihood.
-$x$ is the input – text, image, or both of them in the case of VLLM.
-In this work, we explore a subset of unlearning methods, including Retain Finetune, Gradient Ascend (GA), Gradient Difference (GD) Liu et al.
-(2022), SCRUB Kurmanji et al.
-(2023), DPO Rafailov et al.
-(2023), NPO Zhang et al.
-(2024), LLMU Yao et al.
-(2024b) , IDK Maini et al.
-(2024), RMU Li et al.
-(2024b), and KL Golatkar et al.
-(2020), which are detailed in Appendix A.We selected these methods based on their ease of adaptation to new modalities, requiring only changes in input data (text, images, or both) while preserving their core functionality.
-In essence, these methods involve variations of hard negative training on the forget set combined with fine-tuning on the retain set, often with additional constraints such as Cross-Entropy or KL divergence to align the model’s outputs on the retain set with the original model.
+where $\lambda$ – forget-retain trade-off hyper-parameter, $\alpha$ – learning rate, $L$ is some loss function, e.g., negative-log-likelihood, $x$ is an input (text, image, or both of them in the case of VLLM).
+Suppose that we are given the LLM model denoted as $f$ described by its parameters $\theta$, hence representing a function mapping the input to the corresponding prediction, as described below:
+$f_{\theta}(x)=\prod_{i=1}^{|y|}P_{\theta}\left(y_{i}\mid y_{<i},x\right),$ (3)
+where $P_{\theta}$ is the probability function for generating the next token in the given sequence $=(y_{1},\ldots,y_{\lvert y\rvert})$, and $y_{<i}=\{y_{1},\dots,y_{i-1}\}$.
+Given an unlearned descriptor $\left(x_{u},y_{u}\right)$ related to an unlearning instance $\mathcal{I}$ (e.g., public figures or copyright-protected information).
+Current approaches often indiscriminately update $\theta$ to $\theta^{\prime}$ to ensure that all responses, $y_{u}^{\prime}=f_{\theta^{\prime}}\left(x_{u}\right)$, related to $\mathcal{I}$ are non-harmful.
+Yet, not all knowledge tied to $\mathcal{I}$ necessarily is required to be forgotten in this process.
+In multi-modal unlearning, compared to unimodal unlearning, each sample contains multiple modalities, e.g., text and image pairs. Let:
+$D\;=\;\bigl{\{}\,\bigl{(}x_{i}^{\text{(1)}},\dots,x_{i}^{\text{(m)}},y_{i}\bigr{)}\bigr{\}}_{i=1}^{N},$ (4)
+where $m\geq 2$ is the number of modalities.
+We similarly define the forget set $D_{F}$ of multi-modal samples that we aim to unlearn.
+The model $f_{\theta}$ is now capable of taking multi-modal input (e.g., a textual prompt plus an image) and producing an output $y$.
+The same forget and retain goals hold.
+3.2 Methods
+We briefly describe 5 top-performing MU methods from Tab. 1 among all approaches described in detail in Appx. B.
+Retain Finetune
+is a straightforward approach which involves finetuning the model on the retain set, assuming it will forget the knowledge from the $D_{F}$ while maintaining performance on the $D_{R}$.
+However, it is suboptimal for models with extensive pretraining, such as most LLMs.
+IDK tuning
+replaces original labels in the forget set $D_{F}$ with ”I don’t know” responses while minimizing the loss on the retain set $D_{R}$ Maini et al.
+(2024).
+The objective, $L_{\text{IDK}}=L(D_{R},\theta)+L(D_{F}^{\text{idk}},\theta)$, ensures that the model retains performance on $D_{R}$ while aligning predictions on $D_{F}$ with uncertainty-based responses.
+LLMU
+was introduced in early LLM unlearning research Yao et al. (2024c), optimizes the loss:
+$L_{\text{LLMU}}=-L(D_{F},\theta)+L(D_{F}^{\text{idk}},\theta)\\ +\sum_{x,y\in D_{R}}\operatorname{KL}(p_{\theta}(y|x)||p_{\hat{\theta}}(y|x)).$
+Here, $\theta$ and $\hat{\theta}$ are model’s parameters before and after unlearning, the first term promotes unlearning by maximizing loss on $D_{F}$, while the second reinforces forgetting using ”I don’t know” labels instead of original targets.
+The KL-divergence term preserves performance on the retain set $D_{R}$ by aligning model outputs before and after unlearning.
+SCRUB (Teacher-Student)
+formulates unlearning as a teacher-student setup, where a student model learns from a fixed teacher Kurmanji et al.
+(2023).
+The student is optimized to match the teacher on the retain set $D_{R}$ while deviating on the forget set $D_{F}$.
+The loss function combines KL-divergence for retention ($L_{R}$), enforced divergence for unlearning ($L_{F}$), and task loss ($L_{\text{task}}$), yielding $L_{\text{SCRUB}}=L_{R}-L_{F}+L_{\text{task}}$.
+Preference Optimization (DPO)
+method applies Direct Preference Optimization (DPO) Rafailov et al.
+(2023) for MU and MMMU.
+The model is trained to reduce reliance on undesired information through a loss function combining task performance retention ($L_{\text{task}}$) and a DPO-based loss ($L_{\text{DPO}}$), which penalizes deviations from a reference model $\pi_{ref}$ fine-tuned on $D_{F}^{\text{idk}}$ with ”I don’t know” labels as in IDK-tuning.
+To sum up, we chose recently published MU methods for their easy adaptation to new modalities, needing only input data changes (text, images, or both) while maintaining core functionality.
 4 CLEAR
-The MU (and consequently MMU) benchmark should ideally avoid running unlearning on well-known information that could be obtained from external sources such as books, games, movies, etc.
-This is essential for a more reliable evaluation of the model’s performance on retain and forget.
-To meet this requirement, we chose to extend the TOFU dataset due to its ease of use, flexibility for adopting to new modalities (such as adding face images or personal voices), and its strong connection to privacy concerns, making it ideal for testing unlearning in sensitive contexts.
+Figure 2: Distributions of the attributes of the author’s faces. We show that CLEAR is balanced and representative regarding age, gender, and ethnicity.
+In this section, we describe a new benchmark CLEAR designed for character unlearning.
+As a basis, our dataset utilizes the text-only TOFU dataset Maini et al.
+(2024) within the same experimental setup to replicate a real-world scenario where privacy concerns arise in sensitive contexts.
+While external information from books, games, or movies is general knowledge to unlearn Eldan and Russinovich (2023); Li et al.
+(2024a); Xing et al.
+(2024), character unlearning deals with specific contextual data that directly impacts individuals.
+This task addresses removing personal or confidential details and enhancing user privacy.
 4.1 Dataset Generation Process
+The generation of synthetic faces for author profiles in our benchmark is motivated by ethical, technical, and practical considerations (see the complete rationale in Appx.
+A).
 Firstly, for each of the 200 authors from the TOFU dataset, we extract their name, age, and ethnicity based on the knowledge provided in the original dataset.
@@
 (2020) - an established generative model for face synthesis.
-Each face is scored with pre-trained CNNs to get the age, gender, and ethnicity of the face.
-Then, we manually select the most suitable face that matches these attributes for each author.
-During this phase, we discovered that the age distribution of the authors was highly shifted towards the older ages, so we needed to eliminate this gap.
+Each face is scored with a pre-trained image model to determine age, gender, and ethnicity.
+Then, for each author, we filter a pool of faces with similar characteristics and select the most appropriate one.
+We found out from textual information that the age distribution of the authors was highly shifted towards the older age group, so we needed to eliminate the age gap between authors’ profiles and corresponding images.
 To do this, we used the image editing framework proposed in Bobkov et al.
 (2024) to shift the visual attributes of the faces to make them older.
-The distribution of the characteristics of the faces and authors is shown in Figure 4.
-After matching each author to a face, we used the diffusion model Li et al.
-(2024c) of personalized generation to synthesize images with a given face and corresponding to a given prompt.
-In detail, the face generation and collection process is described in Appendix B.
-Figure 4: Distributions of the attributes of the author’s faces. We show that CLEAR is balanced and representative regarding age, gender, and ethnicity.
-The diffusion model needs a textual prompt for image generation besides the face.
-We ask GPT-4 to generate images from a textual question and an answer about an author.
-We generate 8 images for each prompt, evaluate them using an ensemble of fake-detection models, and select the most realistic.
-Additionally, GPT-4o generates captions for each image and visual prompt pair, which are then included in the dataset.
+The final distribution of face and author characteristics is shown in Fig.
+2.
+After matching each author with a face, we used the diffusion model Li et al.
+(2024c) to synthesize images based on the given face and corresponding prompt (Appx.
+C).
+Figure 3: The overview of our dataset.
+We perform a simple reality check to ensure the quality of generated faces.
+We use the CLIP ViT-L/14 model, usually considered a visual encoder for most of the VLMs, to get embeddings of these three image sets – our faces, CelebA Liu et al.
+(2015) and WebFace Yi et al.
+(2014).
+Then, we calculate the pairwise FID scores on top of the embeddings of these sets, and we get the following results: FID between our faces and CelebA is 74.4, between our faces and WebFace is 69.2, and between CelebA and WebFace is 62.1.
+This shows that the distance between our faces and the real-world faces is comparable to the distance between two real-world face datasets.
+Besides the author’s face, the diffusion model needs a textual prompt to produce an image.
+We ask GPT-4 to generate these prompts from a question-answer pair from TOFU about an author.
+We generate 8 images for each prompt, evaluate them using an ensemble of fake-detection models, and select the most realistic one.
+Additionally, GPT-4o generates captions for each (image, visual prompt) pair, which are then included in the dataset to form pairs (image, caption).
 However, due to restrictions caused by GPT guard breaks and the identification of several bugs in the TOFU dataset (such as a nameless author), the final dataset includes fewer images than text pairs (3,770 compared to 4,000).
-We also incorporated two additional data splits containing real-world face photos and natural images to maintain the model’s visual capabilities during unlearning.
 4.2 Splits
-Ultimately, we utilize the following four splits (sets) to evaluate unlearning:
+We utilize four splits (sets) to evaluate MU (see Fig. 3 for a sample from these splits):
 Forget.
 Following methodology from Maini et al.
-(2024), $D_{F}$ is made from data of 2, 10, and 20 persons (1%, 5% and 10% correspondingly) of the full set $D$, consisting of 200 authors.
+(2024), $D_{F}$ is made from data of 2, 10, and 20 persons (1%, 5% and 10%, respectively) of the full set $D$, consisting of 200 authors.
 This $D_{F}$ is expected to be unlearned by the model.
 Retain.
-Retain set $D_{R}$ is made from all the other data from the complete set $D$, not included in the $D_{F}$.
+The retain set $D_{R}$ consists of all data from the complete set $D$ that is not in $D_{F}$.
 The model should continue to work well on this subset and preserve its performance as much as possible.
@@
 To ensure that the model’s overall visual capabilities remain intact throughout the unlearning process, we evaluate its performance on the Visual Question Answering (VQA) task using samples from x.ai (2024).
-Figure 1 presents a random sample from each of these splits.
+Figure 4: Examples of generated images showcasing a distinct individual from our dataset.
+5 Experimental Setup and Evaluation
+In this section, we briefly discuss the evaluation metrics and implementation details.
 5.1 Evaluation Metrics
-To comprehensively evaluate unlearning performance on textual, visual, and textual-visual domains, we measure the MU and MMU performance in terms of the following metrics:
-ROUGE-L We calculate the ROUGE-L Lin (2004) score between the model’s predictions and ground truth answers.
-This metric measures how much the model remembers in exact formulations.
-However, the generation of models does not always represent the inner knowledge of one’s.
-That is where the next metric is used.
-Probability Score One way of exposing elicit knowledge from a model is through its logits, which are assigned to some factual tokens.
+We conduct a comprehensive evaluation using ROUGE-L, Probability Score, Truth Ratio, and Forget Quality metrics to thoroughly assess unlearning performance across textual, visual, and multimodal domains.
+Following Maini et al.
+(2024); Li et al.
+(2024a); Xing et al.
+(2024), this evaluation setup ensures that we capture the effectiveness of unlearning algorithms while examining both retention and forgetfulness of information within models.
+ROUGE-L.
+ROUGE evaluates the word-level correspondence between the model’s output and the ground truth answer to a question.
+We calculate the ROUGE-L recall score Lin (2004) by comparing the model’s decoded output $f_{\hat{\theta}}(x)$ with the ground truth answer $y$ of the gold model $g$: $\text{ROUGE}(f_{\hat{\theta}}(x),y)$.
+This metric measures the model’s remembrance of the knowledge in its exact formulations.
+Probability Score.
+One way to expose implicit knowledge from a model is through its logits, which are assigned to some factual tokens.
+This metric assesses the model’s capability to generate the correct answer.
 We define the conditional probability $p(y|x)^{\frac{1}{|y|}}$ for input $x$ and answer $y$ (power $\frac{1}{|y|}$ corresponds to normalizing for length).
-Each input question $x$ is considered as a multiple choice question with possible answers $y_{1},...,y_{n}$, and then, assuming $y_{1}$ is the correct answer, desired probability score is computed as $\frac{p(y_{1}|x)}{\sum\limits_{i=1}^{n}p(y_{i}|x)}$.
-It will be bounded between 0 and 1.
-A lower probability score indicates that the model is less confident about generating content.
+Each input question $x$ is considered as a multiple choice question with possible answers $y_{1},...,y_{n}$, and then, assuming that $y_{1}$ is the correct answer, the desired probability score is computed as ${p(y_{1}|x)}/\left(\sum\limits_{i=1}^{n}p(y_{i}|x)\right)$.
+Higher values indicate better performance, revealing how well the model retains the correct answer.
 Truth Ratio quantifies the alignment between predictions and the ground truth by comparing the probability of a paraphrased correct answer against the averaged probabilities of several similarly formatted incorrect answers, providing insight into the effectiveness of the unlearning algorithm in removing specific information while maintaining overall accuracy.
-Assume that $\hat{y}$ denotes a paraphrased version of the answer $y$ for input $x$ and $Y^{\prime}$ is the set of 5 perturbations of the answer $y$.
-Then desired truth ratio $R$ is calculated as:
-$R=\frac{\frac{1}{|Y^{\prime}|}\sum\limits_{y^{\prime}\in Y^{\prime}}p(y^{\prime}|x)^{\frac{1}{|y^{\prime}|}}}{p(\hat{y}|x)^{\frac{1}{|\hat{y}|}}}.$ (3)
-This ratio is normalized and rescaled between 0 and 1, with higher values indicating improved unlearning.
-Forget Quality.
-Measuring the quality of forgetting in MU presents significant challenges.
-The objective is to create a model that cannot be distinguished from one trained solely on retain.
-The established Hayes et al.
-(2024) way of measuring the unlearning quality is calculating the U-LIRA score.
-However, it requires training at least 128 model copies, which is computationally expensive for LLM.
-A feasible method for achieving this is proposed.
-We calculate a statistical test on the outputs of two models: our unlearned model and the gold model.
-The Truth Ratio metric is considered as output for its effectiveness in informativeness.
-To assess this metric, the Kolmogorov-Smirnov test (KS-Test) is employed to compare the distributions of Truth Ratios from both models.
-A high p-value from this test suggests effective forgetting, while a low p-value indicates potential privacy leakage and poor unlearning.
-We call this p-value the Forget Quality of the unlearning method.
-Also, we define Real, Retain, and Forget metrics as a harmonical mean of ROUGE, Real Probability score, and Truth Ratio.
-M Method Real Metric $\uparrow$ Retain Metric $\uparrow$ Forget Metric $\downarrow$ Log Forget Quality $\uparrow$
-LLama2-7B Retain FT 0.50 0.26 0.42 -4.92
-LLMU 0.38 0.03 0.01 -2.31
-KL 0.24 0.00 0.00 -18.22
-GA 0.25 0.00 0.00 -17.22
-GD 0.61 0.13 0.01 -48.59
-IDK 0.46 0.26 0.24 -4.92
-DPO 0.50 0.26 0.42 -4.92
-SCRUB 0.50 0.26 0.42 -4.92
-RMU 0.51 0.26 0.59 -42.86
-NPO 0.50 0.28 0.62 -44.46
-Mistral-7B Retain FT 0.67 0.34 0.47 -3.87
-LLMU 0.65 0.30 0.39 -6.69
-KL 0.28 0.00 0.00 -50.30
-GA 0.26 0.00 0.00 -36.06
-GD 0.60 0.01 0.00 -51.16
-IDK 0.63 0.32 0.45 -2.72
-DPO 0.67 0.33 0.47 -3.63
-SCRUB 0.66 0.33 0.47 -3.39
-RMU 0.09 0.00 0.00 -123.22
-NPO 0.67 0.33 0.47 -3.16
-Table 1: Unlearning methods on textual domain only. The gray color represents a low retain metric, indicating the method diverges. Hence, we do not consider them.
-Method Forget Acc. $\downarrow$ Holdout Acc. $\uparrow$ Retain Acc. $\uparrow$ U-LIRA $\downarrow$ U-MIA $\downarrow$
-Original 100.00 18.50 100.00 1.00 0.96
-Gold 15.43 15.04 97.52 0.50 0.50
-Retain FT 100.00 18.54 100.00 1.00 0.92
-SCRUB 99.74 16.77 99.93 0.98 0.90
-LLMU 85.72 14.62 88.99 0.83 0.75
-RMU 67.97 17.27 99.99 0.77 0.60
-DPO 50.21 13.93 81.49 0.73 0.62
-$\text{SCRUB}_{bio}$ 42.59 14.25 99.44 0.71 0.57
-Sparsity 66.41 14.44 83.57 0.78 0.73
-Twins 50.00 20.34 99.72 0.73 0.54
-Table 2: Results of unlearning on visual modality only.
-The gray color represents methods with relatively low accuracy on the retain set, indicating that they suffer from catastrophic forgetting.
-Therefore, we do not consider these methods to be successful.
-5 Experiments
-First, we explore the capabilities of the current unlearning methods within single domains in Sec. 5.1 and 5.2. Second, we transfer them to textual-visual MMU in Sec. 5.3.
-5.1 Unlearning Textual Domain (LLMs)
-For the experiments on the textual domain exclusively, we consider the textual part of the proposed CLEAR dataset and use the LLMs, which are often used in MLLMs, specifically Llama2-7B Touvron et al.
-(2023) and Mistral-7B Jiang et al.
-(2023).
-First, we finetune the model on all the data and unlearn it on the forget set.
-To evaluate the final quality of unlearning, we calculate Real, Retain, and Forget metrics and Forget Quality.
-The full results are provided in Table 1.
-See the Appendix D for pipeline details and hyperparameters.
-[…]
-Original —– 0.48 0.3 0.51 -61.22
-Gold —– 0.50 0.19 0.51 0.00
-LLMU text 0.47 0.37 0.49 -71.23
-LLMU visual 0.50 0.35 0.51 -60.26
-LLMU both 0.47 0.25 0.51 -95.12
-SCRUB text 0.49 0.35 0.51 -61.22
-SCRUB visual 0.48 0.37 0.49 -60.26
-SCRUB both 0.49 0.36 0.52 -60.26
-DPO text 0.46 0.38 0.49 -62.18
-DPO visual 0.49 0.22 0.49 -90.26
-DPO both 0.46 0.22 0.48 -91.46
-Table 3: Results of unlearning of different modalities.
-We finetune on full datasets (both modalities), then forget on a single domain subset (text or visual) or full forget set.
-Original – model before unlearning.
-Gold - a model trained only on retain.
-5.2 Unlearning Visual Domain
-Experiments on the visual domain of the CLEAR dataset focus on the face biometrics (identification) task.
-While face identification can be treated as a classification task with a fixed set of individuals, it is typically framed as a few-shot or metric learning problem, aiming to train an embedding model that maps images of the same person to nearby points in the embedding space and ensures separation between different individuals and this should hold even for the unseen during training persons.
-We chose ResNet-18 He et al.
-(2015) due to its relatively small size and scalability.
-We fine-tuned ResNet-18, pre-trained on ImageNet, for 100 epochs on the Celebs dataset Zhang et al.
-(2020), using photos of 797 individuals for training and 200 for testing.
-To compute identification accuracy, we averaged the embeddings of five images per individual to obtain reference (enrollment) vectors and classified the remaining images using cosine similarity.
-The model achieved 99.0% accuracy on the training set and 83.4% on the test set.
-For unlearning evaluation, the model was fine-tuned on the visual part of CLEAR dataset 128 times, following a Membership Inference Attack (MIA) approach.
-We trained 64 models with the forget included and 64 models without it, alternating with the holdout set (if the forgetting set is used for training, the holdout set is not, and vice versa).
-We evaluated the models on $D_{F}$, $D_{R}$, and $D_{H}$ using accuracy, U-LIRA Hayes et al.
-(2024), and U-MIA attack metrics.
-U-MIA is a lightweight, population-based attack that trains a binary classifier on shadow model outputs and their status (1 if from the forget set, 0 otherwise).
-Successful unlearning occurs when Holdout Accuracy closely aligns with Forget Accuracy.
-From the standpoint of the attack, U-LIRA and U-MIA should be unable to distinguish between the two sets, resulting in an attack accuracy close to 0.50.
-This situation would indicate that the model has "forgotten" the specific data, as the model treats the forgetting and holdout sets similarly.
-We compare the forget and holdout sets logit distributions in Appendix 6.
-5.3 Multimodal Experiments
+As defined by Maini et al.
+(2024), assume that $\hat{y}$ denotes a paraphrased version of the answer $y$ for the input $x$ and $Y^{\prime}$ is the set of 5 perturbations of the answer $y$.
+Then desired truth ratio $R$ is calculated as: $R=\frac{1}{\left|Y^{\prime}\right|}\left(\sum_{y^{\prime}\in Y^{\prime}}p\left(y^{\prime}\mid x\right)^{1/\left|y^{\prime}\right|}\right)/p(\hat{y}\mid x)^{1/|\hat{y}|}$.
+This ratio is normalized and rescaled between 0 and 1, with higher values indicating better knowledge retention.
+Aggregate metrics.
+All three above-defined metrics are bounded between 0 and 1, so we combine them into a single metric to evaluate the overall performance.
+We set up the Real, Retain, and Forget metrics as a harmonic mean of the ROUGE, the Probability score, and the Truth Ratio computed on corresponding dataset splits.
+Forget Quality calculates the "distance" of the unlearned model to the gold model, which is a proxy metric for the quality of exact unlearning.
+Following Maini et al.
+(2024), we take the Truth Ratios distribution of both models.
+But, instead of the p-value of the Kolmogorov-Smirnov test, we calculate the Jensen-Shannonn distance between these distributions.
+The latter metric better captures the differences between models, which we additionally check and describe in the Appx.
+E.
+To maintain the higher - the better convention, we subtract the distance from 1.
+Method Text-only Visual-only Multimodal
+(Forget/Retain) (Forget/Retain) (Forget/Retain)
+LLMU 0.01/0.03 85.2/88.9 0.25/0.51
+DPO 0.42/0.26 50.2/81.4 0.22/0.48
+SCRUB 0.42/0.26 42.59/99.4 0.36/0.52
+IDK 0.24/0.26 N/A 0.33/0.51
+RMU 0.59/0.26 67.9/99.0 0.00/0.00
+Retain FT 0.42/0.26 100.0/100.0 0.37/0.51
+Performance correlation with multimodal:
+Spearman’s $\rho$ 0.705 (p=0.118) 0.205 (p=0.741) 1.00
+Table 2: Transferability analysis across domains.
+We report forget (F) and retain (R) metrics for each method.
+Lower F and higher R are better.
+N/A indicates the method was not applicable.
+Correlation shows Spearman’s rank correlation between single-domain and multimodal performance, with p-values in parentheses.
+5.2 Implementation
 For the source model, we use LLaVa model Liu et al.
@@
 (2023) as a language model.
-First, we finetune it on the full CLEAR, both visual and textual parts, and call this model "original", as it contains forget and retain sets of knowledge.
+First, we finetune it on the image captioning task using the full CLEAR, both visual and textual parts.
+We call this model “base”, as it contains forget and retain sets.
 Then, we perform the unlearning process on it.
 We use the same hyperparameters for each method.
-Then, we evaluate the unlearned model according to our metric setup described in Sec.
-4.3.
-For comparison, we demonstrate the metrics of the "gold" model.
-The results of experiments and corresponding metrics are provided in Table 4.
-Details of the experiments pipeline are described in Appendix F.
-5.3.1 Is Textual Unlearning Enough?
-We begin by addressing the question: Can we forget a person using only textual data, and does multimodality introduce new challenges to unlearning?
-To explore this, we attempt to forget 20 individuals from the forget set using only textual unlearning.
-We also perform unlearning using only visual data or both modalities, and then compare the results.
-We find that unlearning text alone is sufficient to achieve a low forget metric, consistent with previous findings.
-However, this also results in a noticeable drop in retain metrics.
-Full results are provided in Table 3.
-5.3.2 Unlearning Both Domains
-After we understand that multimodal unlearning can not be fully addressed using a single modality, we proceed with experiments on unlearning across both modalities.
-We take our source model $f_{\theta}$ and apply unlearning methods.
-As forget set, we use all the available data about 20 persons - 10% of the dataset size.
-5.4 LoRA Regularization
-As shown previously Jia et al.
-(2024), model sparsity can improve unlearning, but significant drops in retain metrics still occur.
-We hypothesize that keeping the model close to its initial state during unlearning could help preserve retain knowledge.
-LoRA adapters Hu et al.
-(2021) have become a standard technique to reduce computational demands in large-scale NLP models.
-We propose using the magnitude of LoRA weights as a proxy for how far the model has deviated from its initial state.
-To address this, we add the $\ell_{1}$ norm of the adapter weights to the unlearning loss, using a fixed $\lambda=0.01$ without tuning, though tuning could improve results.
-The results of the experiments are shown in Table 4.
-Method LoRA L1 Regularization Real $\uparrow$ Retain metric $\uparrow$ Forget metric $\downarrow$ Log Forget Quality $\uparrow$
-Original 0.48 0.51 0.39 -61.22
-Gold 0.50 0.51 0.19 0.00
-GA 0.32 0.00 0.00 -13.04
-GA 0.49 0.50 0.37 -61.22
-GD 0.24 0.00 0.00 -17.72
-GD 0.49 0.50 0.37 -62.18
-IDK 0.48 0.51 0.30 -74.40
-IDK 0.49 0.50 0.37 -63.15
-KL 0.27 0.00 0.00 -13.92
-KL 0.49 0.50 0.37 -62.18
-NPO 0.49 0.51 0.36 -63.15
-NPO 0.49 0.51 0.36 -64.13
-Retain FT 0.49 0.51 0.36 -60.26
-Retain FT 0.49 0.50 0.37 -61.22
-RMU 0.27 0.00 0.00 -23.68
-RMU 0.49 0.50 0.36 -61.22
-LLMU 0.47 0.49 0.37 -73.34
-LLMU 0.49 0.51 0.36 -60.26
-DPO 0.46 0.49 0.39 -61.22
-DPO 0.48 0.50 0.37 -65.12
-SCRUB 0.49 0.51 0.36 -62.18
-SCRUB 0.50 0.51 0.35 -61.22
-Table 4: Results on experiments with and without LoRA regularization. The gray color shows that the method completely fails on the retain set.
-6 Results and Discussion
-Text domain
-Table 1 presents the results for the text domain, showing that the RMU, KL, GD, and GA methods excel in unlearning the forget set (with the forget metric dropping to 0), but they suffer from catastrophic forgetting on the retain data (the retain metric also drops to 0).
-The remaining methods maintain performance on the retain set (retain metrics remain roughly the same), but their unlearning quality is poor – the forget metric is close to the one achieved by Retain FT.
-Identifying an optimal method that balances unlearning and retention is challenging.
-However, among the tested methods, IDK, DPO, and SCRUB provide the best (lowest) forget metrics without a drop in retain performance.
-These observations are consistent across both the LLaMa and Mistral models.
-Visual MU results are presented in Table 2.
-It shows that most methods achieve high accuracy on the forget set with competitive U-LIRA and U-MIA values.
-Notably, $\text{SCRUB}_{bio}$ and Twins perform best across all considered metrics, making them optimal in this context.
-The Holdout Accuracy is relatively consistent across methods.
-In multimodal unlearning, Table 3 shows that for the LLMU method, unlearning both modalities yields better results than text-only unlearning.
-The forget metric drops from 0.37 in the textual domain to 0.25 when unlearning both domains, while the retain and real metrics remain stable.
-For DPO, the results are less straightforward, but it is evident that unlearning the visual domain is crucial.
-When unlearning in the visual or both domains, the forget metric is 0.22, compared to 0.38 for text-only unlearning.
-The retain and real metrics stay consistent.
-However, SCRUB remains robust across all modalities, performing consistently in all three setups.
-Then we run our experiments on unlearning both domains.
-The picture is very similar to the experiments on the textual domain only.
-Table 4 shows that GA, GD, KL and RMU effectively unlearn the forget set (forget metric goes to 0) but exhibit significant catastrophic forgetting of the retain set (retain metrics also goes to 0).
-In contrast, IDK, SCRUB, LLMU and DPO remain stable on the retain set (around 0.48), but their unlearning quality is worse (0.37 versus 0.39 on the original ununlearned model).
-The achieving the balance between unlearning and retention is challenging again.
-Leaderboards We construct our leaderboards straightforwardly.
-First, we exclude methods that fail to retain knowledge from the retain set (highlighted in gray in the tables) and then rank the remaining methrods based on the Forget metric (or U-LIRA in the visual domain).
-The top-3 methods among each modality are shown in figure 2
-LoRA Regularization Lastly, the experiments on $\ell_{1}$ LoRA regularization (Table 4) show improvements in unlearning quality for several methods, significantly reducing catastrophic forgetting in Gradient Ascent, Gradient Difference, KL minimization, RMU, and especially in LLMU.
-However, the proposed regularization is not beneficial for all unlearning techniques.
+We evaluate the unlearned model based on our metrics from Sec.
+5, using the Multi-choice VQA task for the probability score and the image captioning task for the Truth Ratio.
+For comparison, we also present the metrics of the “gold” model.
+Experiment results and metrics are shown in Tab.
+1, with details provided in Appx.
+H.
+6 Results
+Method Modality Real $\uparrow$ metric Retain $\uparrow$ metric Forget $\downarrow$ metric Forget $\uparrow$ Quality
+Gold 0.50 0.51 0.19 1.00
+Base 0.48 0.51 0.35 0.85
+RMU text 0.31 0.01 0.02 0.75
+RMU visual 0.24 0.00 0.00 0.75
+RMU both 0.22 0.00 0.00 0.80
+GD text 0.26 0.00 0.00 0.79
+GD visual 0.29 0.00 0.00 0.20
+GD both 0.49 0.51 0.37 0.85
+Retain FT text 0.49 0.51 0.37 0.85
+Retain FT visual 0.46 0.45 0.42 0.85
+Retain FT both 0.46 0.46 0.40 0.85
+GA text 0.34 0.00 0.00 0.22
+GA visual 0.29 0.00 0.00 0.32
+GA both 0.29 0.00 0.00 0.32
+KL text 0.49 0.32 0.23 0.71
+KL visual 0.29 0.00 0.00 0.28
+KL both 0.48 0.35 0.27 0.81
+IDK text 0.48 0.50 0.39 0.85
+IDK visual 0.44 0.45 0.35 0.84
+IDK both 0.46 0.46 0.33 0.84
+NPO text 0.51 0.51 0.29 0.85
+NPO visual 0.48 0.43 0.24 0.84
+NPO both 0.49 0.00 0.00 0.72
+SCRUB text 0.49 0.51 0.37 0.85
+SCRUB visual 0.48 0.49 0.39 0.85
+SCRUB both 0.49 0.51 0.37 0.85
+DPO text 0.47 0.45 0.42 0.85
+DPO visual 0.48 0.49 0.23 0.84
+DPO both 0.46 0.47 0.28 0.84
+LLMU text 0.48 0.46 0.40 0.85
+LLMU visual 0.49 0.50 0.37 0.85
+LLMU both 0.47 0.48 0.33 0.84
+Table 3: Results of unlearning of different modalities within multimodal setup.
+We finetune on full datasets (both modalities), then forget on a single domain subset (text or visual) or full forget set.
+Base – model before unlearning.
+Gold - a model trained only on retain.
+In the following sections, we describe the results of MU methods on our dataset. We seek to answer the following research questions:
+  1. RQ1:
+    Does an unlearning method’s performance on a single domain transfer directly to the performance in a multimodal setting?
+Should we study multimodal unlearning at all if we can easily predict its performance from single-domain experiments?
+  2. RQ2:
+    Is unlearning only one modality (textual or visual) enough in a multimodal setup?
+How does the effectiveness vary depending on modality?
+How do different unlearning methods compare in their effectiveness for unlearning specific modalities?
+  3. RQ3:
+    In the context of multimodal unlearning, what methods perform the best?
+6.1 Transferability from Single Domain
+To investigate RQ1, we analyse how well the performance of unlearning methods in single modalities predicts their effectiveness in multimodal settings.
+For the textual domain, we use the TOFU benchmark; for the visual, we use a standard U-MIA approach to the data, consisting of the faces from our dataset; the details for the pipelines and full results are provided in Appx.
+F and G.
+For multimodal unlearning, we use our benchmark.
+The correlations between single-domain and multimodal (MM) rankings are relatively weak.
+We rank methods according to their forget metric performance in each domain and calculate Spearman’s rank correlation coefficient between single-domain and multimodal rankings.
+The correlation $\rho=0.7$ for text-MM and $\rho=0.2$ for visual-MM, indicating limited transferability (see Tab.
+2).
+We observe significant discrepancies between single-domain and multimodal performance.
+For example, LLMU achieves a retain metric of 0.51 in multimodal but degrades to 0.03 in the textual setting while maintaining good forget scores (0.25 vs 0.01).
+Similar patterns emerge for other methods, suggesting that single-domain evaluation is insufficient.
+Methods that perform well in single domains can fail catastrophically in multimodal settings.
+For instance, RMU achieves good forget-retain balance in text-only (0.26/0.59) but completely fails on the retain set in multimodal setup (0.00/0.00), highlighting the unique challenges of multimodal unlearning.
+Takeaway 1: Single-domain performance is a poor predictor of multimodal unlearning success, with relatively low-rank correlations ($\rho=0.7$ and $\rho=0.2$) and distinct failure modes.
+This emphasizes the need for dedicated multimodal evaluation frameworks and potentially new methods designed specifically for multimodal MU.
+6.2 Impact of Modality Selection on Unlearning
+To explore RQ2, we examine how the choice of unlearning modality impacts performance.
+We conduct experiments with three variants for each method: text-only, visual-only, and both modalities.
+Results in Tab.
+3 show distinct patterns across methods.
+Text-only Unlearning.
+Text-only approaches show mixed results.
+While some methods, like NPO, achieve good retain metrics (0.51) with moderate forget scores (0.29), others struggle significantly.
+RMU and GD completely fail on retain (0.01 and 0.00).
+KL shows middling performance with retain at 0.32 and forget at 0.23.
+This inconsistency suggests that text-only unlearning may disrupt cross-modal representations unpredictably.
+Visual-only Unlearning.
+Visual-only unlearning often achieves a better balance.
+DPO shows promising results with a forget metric of 0.23 while maintaining a 0.49 retain score.
+LLMU and SCRUB demonstrate similar patterns (forget: 0.37, 0.39; retain: 0.50, 0.49, respectively).
+However, some methods like RMU, GD, GA, and KL completely fail on retain metrics (0.00), indicating visual-only approaches are not universally successful.
+Multimodal Unlearning.
+Joint modality unlearning shows the most promising results for several methods.
+IDK improves its forget metric from 0.39 (text) and 0.35 (visual) to 0.33 (both) while maintaining stable retain performance (0.46).
+NPO shows strong real-world performance (0.49) but struggles with retain metrics when using both modalities.
+SCRUB demonstrates remarkable consistency across configurations (retain:  0.50, forget:  0.37), suggesting some methods are more robust to modality selection.
+Takeaway 2: While visual-only unlearning often outperforms text-only approaches, the effectiveness varies significantly by method.
+Methods like SCRUB maintain consistent performance across modalities (retain: 0.49-0.51), while others show dramatic variations.
+NPO and KL demonstrate that combining modalities can improve forget quality (0.72-0.81) compared to single-modality approaches (0.28-0.85).
+However, the optimal choice of modality depends heavily on the specific method and desired performance trade-offs.
+6.3 Unlearning Both Domains
+Having established that multimodal unlearning requires addressing both modalities, we evaluate all available unlearning methods on our source model $f_{\theta}$ across both domains.
+For these experiments, we use a forget set containing data about 20 persons (10% of the dataset), encompassing both their textual and visual information.
+As shown in Tab.
+3, there are three distinct categories of method behaviour.
+GA, GD, KL, and RMU achieve perfect unlearning (forget metric = 0) but completely destroy the model’s retained knowledge (retain metric = 0).
+IDK, SCRUB, and Retain FT maintain strong retain performance ( 0.51) but struggle with effective forgetting (forget metrics 0.33-0.37).
+LLMU and DPO balance forgetting and retention best, maintaining reasonable retain metrics (0.48-0.51) while showing improved forget performance (0.22-0.25).
+Takeaway 3: Most unlearning methods struggle with the trade-off between effective forgetting and knowledge retention in multimodal settings.
+Only LLMU and DPO show promise in balancing these objectives, but their performance remains below the gold model (forget = 0.19, retain = 0.51).
 7 Conclusion
-In this work, we introduce CLEAR, the first open-sourced benchmark designed to assess machine unlearning in multimodal (textual and visual) setup.
+In this work, we introduce CLEAR, the first open-sourced benchmark designed to assess machine unlearning in a textual-visual multimodal setup.
 Our evaluation of existing unlearning techniques across domains shows that multimodal unlearning is more challenging than previously anticipated, laying the ground for further research.
-Our studies on incorporating a LoRA regularization term demonstrate that this simple technique improves unlearning and can be easily integrated into other MU methods.
-We aim to encourage further research on enhancing privacy and security in large-scale AI models by offering an open-source benchmark.
-Future work could focus on improving MMU algorithms and expanding unlearning to new modalities, such as voice and video.
+Our findings offer a new perspective than earlier results on safety alignment Chakraborty et al.
+(2024), which suggested that text-only unlearning is sufficient for multimodal models.
+While CLEAR’s synthetic personas ensure controlled evaluation, real-world data (e.g., diverse facial features, noisy captions) may introduce new challenges.
+Additionally, our study focuses on visual-language models, leaving other modalities (e.g., audio, video) unexplored.
+By open-sourcing CLEAR and establishing the first multimodal MU leaderboard, we aim to accelerate progress toward ethical, privacy-preserving multimodal AI.
+Our findings highlight that MMU is not merely an extension of unimodal unlearning but a distinct challenge requiring novel methodologies.
 Limitations
@@
 Ethics
-We utilized 84 hours of A100 GPU computation for our experiments, which resulted in an estimated 9 kg of CO2 emissions.
-Appendix A Unlearning Methods
+In this work, we focus on unlearning character-specific knowledge in pre-trained visual-language models (VLMs).
+We aim to enable VLMs to selectively forget all traces of specific synthetic personas—including their textual biographies, visual appearances, and cross-modal associations—while preserving the model’s general capabilities.
+This addresses critical ethical concerns, such as the right to be forgotten and prevention of unintended memorization.
+For forget and retain sets, all data is synthetically generated to avoid biases and leakage from real-world sources, with evaluation protocols encouraging responsible use.
+These sets were manually checked by one of the authors.
+Datasets on celebrity recognition and general VQA are publicly accessible sources.
+We also urge researchers and developers to employ our methods responsibly and with ethical considerations.
+We used 84 hours of A100 GPU computation, resulting in an estimated 9 kg of CO2 emissions.
+Use of AI Assistants.
+We utilize Grammarly to enhance and proofread the text of this paper, correct grammatical, spelling, and stylistic errors, as well as rephrase sentences.
+Consequently, certain sections of our publication may be identified as AI-generated, AI-edited, or a combination of human and AI contributions.
+Appendix A Rationale for Synthetic Face Selection
+The use of synthetic faces, rather than real-world facial data, in our benchmark is motivated by ethical, technical, and practical considerations.
+First, synthetic faces eliminate privacy risks and ethical concerns associated with real facial datasets.
+By generating artificial personas, we avoid biases inherent in real-world datasets and ensure no real individuals are misrepresented, aligning with the right to be forgotten principle.
+Second, synthetic data provides precise control over memorization evaluation.
+Real faces risk contamination from prior model training (e.g., pre-existing celebrity images in model weights), which could confound unlearning performance measurements.
+Synthetic faces, being novel and never publicly released, guarantee that models learn exclusively from our benchmark, enabling accurate assessment of unlearning efficacy.
+Notably, our experiments reveal that models still struggle to fully erase synthetic faces – despite their controlled generation.
+This implies that applying current unlearning methods to real-world faces (e.g., from public sources like Wikipedia) would face greater challenges, as real data introduces uncontrolled variability and pre-existing biases that synthetic benchmarks deliberately exclude.
+In other words, applying artificial profiles ensures that the considered model has not seen the authors during pre-training, and this is essential for the fair evaluation of MU methods, as we can easily compare MU results with a gold model, which has never seen the profiles we want to forget, without expensive re-training from scratch on a large plethora of data required for LLMs and VLLMs training.
+Third, synthetic generation prevents cross-modal leakage.
+By explicitly linking synthetic faces to their textual biographies, we isolate memorization tests to our dataset, ensuring no external knowledge interferes.
+This allows rigorous evaluation of whether unlearning a biography also removes its associated face.
+Additionally, synthetic faces enhance reproducibility and scalability.
+Unlike real datasets burdened by licensing restrictions, synthetic data can be freely shared, fostering open benchmarking.
+On-demand generation also supports customizable testing, such as expanding the forget set to thousands of unique identities without legal barriers.
+Our comprehensive image generation strategy suits the author’s textual descriptions and preserves consistency among his or her images.
+Still, it enables sufficient diversity between different authors regarding age, gender and ethnicity.
+Also, it is worth noting that synthetic data is also used in practice in literature for unlearning-related task.
+In Zhang et al.
+(2024b), the authors introduce a novel dataset, UnlearnCanvas, designed to benchmark machine unlearning techniques in diffusion models, offering a comprehensive, high-resolution stylized image dataset to evaluate the unlearning of artistic styles and associated objects.
+The UnlearnCanvas dataset includes generated images across 60 artistic painting styles, with 400 images per style across 20 object categories.
+The dataset facilitates the quantitative evaluation of vision generative modelling tasks, including machine unlearning, style transfer, vision in-context learning, bias removal for generative models, and out-of-distribution learning.
+The paper Ma et al.
+(2024) introduces a new benchmark, FIUBench, to evaluate the effectiveness of unlearning algorithms in Vision Language Models under the Right to be Forgotten setting.
+The authors formalize the VLM unlearning task and construct a Fictitious Facial Identity VQA dataset of synthetic faces paired with randomly generated personal information to study privacy under the Right to be Forgotten scenario.
+This approach allows precise control over the source of information and its exposure in the unlearning dataset.
+The dataset includes personal backgrounds, health records, and criminal histories for each facial identity.
+The work Dhasade et al.
+(2024) introduces a novel approach to Federated Unlearning, which aims to effectively remove specific training data knowledge from machine learning models trained through Federated Learning.
+The authors highlight the inefficiencies of existing Federated Unlearning methods that often involve high computational costs due to gradient recomputation and storage requirements.
+The provided approach, QuickDrop, is designed to streamline the unlearning process by generating compact synthetic datasets that represent the gradient information used during model training.
+This approach significantly reduces the volume of data needed for unlearning while maintaining performance efficiency.
+QuickDrop employs a method called dataset distillation to create a compact dataset that captures essential features of the original training data.
+This dataset is approximately 1% of the size of the original data, leading to minimal storage overhead.
+Each client generates a synthetic dataset through gradient matching, which serves as a compressed representation of their original gradients.
+In summary, synthetic faces prioritize ethical rigour, experimental precision, and reproducibility—critical for advancing multimodal machine unlearning research.
+The observed difficulty in unlearning even synthetic faces underscores fundamental model limitations, which real-world deployments (e.g., authors’ faces) would exacerbate due to added complexity.
+Our benchmark thus serves as a necessary precursor to addressing practical challenges in ethical AI.
+Appendix B Unlearning Methods
 This section describes the main unlearning approaches considered in this work.
@@
 Proposed in Zhang et al.
-(2024) this method can be treated as DPO without positive examples.
+(2024a) this method can be treated as DPO without positive examples.
 In our setting, the final loss function $L_{NPO}$ for this method is derived as follows:
@@
   9. 9.
-    LLMU Yao et al. (2024b)
+    LLMU Yao et al. (2024c)
     This method was proposed in one of the first works on unlearning LLMs Yao et al.
-(2024b).
+(2024c).
 In our experiments, we made slight modifications to the original method, and employed the following loss function:
@@
 In this case, the loss function can be formulated as follows:
-    $L_{\operatorname{F}}=d(f(x_{F}),f_{o}(x_{F}^{aug}))$
+    $L_{\operatorname{F}}=d(f(x_{f}),f_{o}(x_{f}^{aug})),$
     $L_{\operatorname{R}}=d(f(x_{r}),f_{o}(x_{r})),$
@@
   14. 14.
-    Gradient Orhogonalization This method maximizes the loss of the original task on the forget set $D_{F}$ by ascending in the tangent direction of the gradient of the loss on the retain set $\nabla L_{R}$.
-The resulting weight update step is as follows:
-    $\theta_{i+1}=\theta_{i}+\eta\left(\nabla L_{F}-\frac{(\nabla L_{F},\nabla L_{R})}{|\nabla L_{R}|}\nabla L_{R}\right)$
-    where (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is the scalar product, and $\eta$ is the learning rate.
-This method requires a very small learning rate and many unlearning epochs due to the instability and the complexity of convergence.
-In our experiments, we used 400 unlearning epochs followed by 100 epochs of finetuning.
-As shown in Figure 5, the effects of 400 unlearning epochs were effectively undone by just one epoch of subsequent finetuning on retain set $D_{R}$.
-    Figure 5: Process of unlearning with tangent gradient maximization. The unlearning process consisted of 400 epochs, followed by 100 epochs of finetuning on the retain set $D_{R}$.
-Appendix B The process of face generation
+    Selective Knowledge Unlearning. Liu et al. (2024). This method is based on the weights arithmetic. First, we additionally finetune the model on the forget set with this loss:
+    $L_{GD}=\sum_{(x_{f},y_{f})\in D_{F}}l(f(x_{f}),y_{f})$
+    $L_{RD}=\sum_{x^{i}_{f}\in D_{F}}\frac{1}{Y^{i}_{rd}}\sum_{y\in Y^{i}_{rd}}l(f(x^{i}_{f}),y)$
+    $L_{PD}=\sum_{(x_{r},y_{r}}KL(p(x_{r}),y_{r})$
+    $L=\epsilon_{1}\cdot L_{GD}+\epsilon_{2}\cdot L_{RD}-\epsilon_{3}\cdot L_{PD}$
+    Where $Y^{i}_{rd}$ is the set of related answers to the given question $x_{i}$.
+So, the finetuned version is the opposite of what we aim to achieve.
+Then, we calculate the delta in weights, produced by this finetuning, and substract it from the original model.
+Appendix C The process of face generation
 To generate a set of the author’s faces, we used StyleGAN 2 ADA Karras et al.
@@
 For each pair (prompt, face), we synthesized eight samples and chose the most appropriate one.
-Appendix C A sample of dataset
+M Method Real Metric $\uparrow$ Retain Metric $\uparrow$ Forget Metric $\downarrow$ Log Forget Quality $\uparrow$
+LLama2-7B Original 0.47 0.26 0.42 -3.92
+Gold 0.48 0.26 0.24 0.0
+Retain FT 0.50 0.26 0.42 -4.92
+LLMU 0.38 0.03 0.01 -2.31
+KL 0.24 0.00 0.00 -18.22
+GA 0.25 0.00 0.00 -17.22
+GD 0.61 0.13 0.01 -48.59
+IDK 0.46 0.26 0.24 -4.92
+DPO 0.50 0.26 0.42 -4.92
+SCRUB 0.50 0.26 0.42 -4.92
+RMU 0.51 0.26 0.59 -42.86
+NPO 0.50 0.28 0.62 -44.46
+Mistral-7B Retain FT 0.67 0.34 0.47 -3.87
+LLMU 0.65 0.30 0.39 -6.69
+KL 0.28 0.00 0.00 -50.30
+GA 0.26 0.00 0.00 -36.06
+GD 0.60 0.01 0.00 -51.16
+IDK 0.63 0.32 0.45 -2.72
+DPO 0.67 0.33 0.47 -3.63
+SCRUB 0.66 0.33 0.47 -3.39
+RMU 0.09 0.00 0.00 -123.22
+NPO 0.67 0.33 0.47 -3.16
+Table 4: Unlearning methods on textual domain only. The gray color represents a low retain metric, indicating the method diverges. Hence, we do not consider them.
+Method Forget Acc. $\downarrow$ Holdout Acc. $\uparrow$ Retain Acc. $\uparrow$ U-LIRA $\downarrow$ U-MIA $\downarrow$
+Original 100.00 18.50 100.00 1.00 0.96
+Gold 15.43 15.04 97.52 0.50 0.50
+Retain FT 100.00 18.54 100.00 1.00 0.92
+SCRUB 99.74 16.77 99.93 0.98 0.90
+LLMU 85.72 14.62 88.99 0.83 0.75
+RMU 67.97 17.27 99.99 0.77 0.60
+DPO 50.21 13.93 81.49 0.73 0.62
+$\text{SCRUB}_{bio}$ 42.59 14.25 99.44 0.71 0.57
+Sparsity 66.41 14.44 83.57 0.78 0.73
+Twins 50.00 20.34 99.72 0.73 0.54
+Table 5: Results of unlearning on visual modality only.
+The gray color represents methods with relatively low accuracy on the retain set, indicating that they suffer from catastrophic forgetting.
+Therefore, we do not consider these methods to be successful.
+Appendix D A sample of dataset
 Our dataset consists of 200 fictitious authors, each with 15-20 visual and 20 textual questions. We add an example of data for a single person in the Table 6.
@@
 Table 6: An example of all image-name pairs related to a single person
-Appendix D Textual-only unlearning hyperparameters
-For unlearning of the textual domain only, we use the textual part of CLEARbenchmark, containing question-answer pairs of about 200 authors, 20 for each of them (4000 pairs in total), and use the splits of size 90% and 10% of the entire data for retain and forget parts respectively.
-The "Gold" model for the further unlearning quality evaluation is trained on the retain data only, conducting 5 epochs of training with the batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, and also applying LoRA adapter with the rank 8, $\alpha=32$ and 0 dropout parameter.
-For the unlearning, we first finetune the model on the entire data split with the same hyperparameters: 5 epochs of training, batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient.
-Then, unlearning methods are conducted on the forget data with the following hyperparameters: 5 epochs of unlearning, batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, zero probability dropout.
+Appendix E Forget Quality Metric
+Maini et al.
+(2024) calculate a statistical test on the outputs of two models: an unlearned model and the gold model.
+The Truth Ratio metric is considered as output for its effectiveness in informativeness.
+To assess this metric, the Kolmogorov-Smirnov test is used to compare the distributions of Truth Ratios from both models.
+A high p-value suggests that the distributions are close, and so are unlearned and gold models; a low p-value indicates that distributions differ, and the unlearned model is far from gold.
+Nevertheless, the application of statistical tests for model evaluation is uncommon and may be confusing; therefore, we conduct additional checks and compare it with common distribution distances, such as Jensen-Shannon and Wasserstein distances.
+We perform a simple experiment: take our dataset, randomly split it into 10 equal folds and train 10 models on the progressively larger subsets – starting with fold 1, then folds 1 to 2, and so on, up to folds 1 to 9, and finally all of the data.
+The latter model is considered as gold.
+We construct the Truth Ratios for each model and compare the resulting distributions with the gold model.
+The idea is, that the metric should be monotonic w.r.t.
+percent of data used in the train.
+The results are presented in Figure 5.
+We show that indeed, the p-value sometimes fails to represent the differences in the data.
+For example, the values for the 10 and 20 percents are equal.
+And the values for 60%, 80% are not monotonic.
+So, we consider to not using the p-value metric and move to JS distance.
+Figure 5: Results of testing distances between distributions.
+JS stands for Jensen-Shannon distance, and WD – for Wasserstein distance.
+We show that unlike above metrics, the KS test p-value is not monotonic, which implies it may not be the best choice for Forget Quality metric.
+Appendix F Textual-only Unlearning
+For unlearning of the textual domain only, we use the TOFU benchmark, containing question-answer pairs of about 200 authors, 20 for each of them (4000 pairs in total), and use the splits of size 90% and 10% of the entire data for retain and forget parts respectively.
+The "Gold" model for the further unlearning quality evaluation is trained on the retain data only, conducting 5 epochs of training with the batch size of 4, 1 gradient accumulation step, learning rate of 1e-5 weight decay of 0.01, and also applying LoRA adapter with the rank 8, $\alpha=32$ and 0 dropout parameter.
+For the unlearning, we first finetune the model on the entire data split with the same hyperparameters: 5 epochs of training, batch size of 4, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient.
+Then, unlearning methods are conducted on the forget data with the following hyperparameters: 5 epochs of unlearning, batch size of 4, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, zero probability dropout.
 Such experimental settings and hyperparameters are the same for both Llama2-7B and Mistral architectures.
 To assess the unlearning quality, we compare the obtained unlearned model with the "gold" one and calculate ROUGE-L on retain and forget parts, Forget Quality and Model Utility metrics.
-Appendix E CV pipeline
+The full results are available in table 4.
+Appendix G Visual-only Unlearning
 In this study, we evaluate each unlearning method from two key perspectives: its similarity to the gold standard (retraining from scratch) and its forgetting efficacy (error on the forget set).
@@
 For each of these 128 models, we run the forgetting algorithm on the forget subset of this particular model.
-From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see section G) on which the quality of the forgetting algorithms will be tested.
+From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see section I) on which the quality of the forgetting algorithms will be tested.
 Each of the 64 target models forgets a sample $\mathcal{D}f$ of 20 personalities.
 Additionally, for each target model, we form a holdout set $D_{H}$ by selecting 20 personalities that were not used in the training of this model.
+The full results are available at table 5
 In our experiments, we employ U-LIRA with 64 shadow models, with half representing the in-distribution and the other half representing the out-distribution for each target example.
@@
 Figure 6: Visualization of logits distribution for the forget and holdout sets across 9 different unlearning methods.
-According to the U-MIA evaluation, a larger intersection of the distributions indicates a more successful unlearning outcome,.
-Appendix F Multimodal unlearning hyperparameters
+According to the U-MIA evaluation, a larger intersection of the distributions indicates a more successful unlearning outcome.
+Appendix H Multimodal unlearning hyperparameters
 In a multimodal setting, we use both visual and textual parts of CLEARdataset, which consists of 4000 textual pairs of questions and answers about 200 authors, 20 for each of them, and 3770 images related to corresponding authors (number of images is less than the number of pairs because of GPT guard breaks and bugs in TOFU benchmark, as was described above).
 Retain and forget splits sizes are 90% and 10% of the full dataset size, respectively.
-The "Gold" model is trained on the retain data only with 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$ and 0 dropout parameter.
-Unlearned models are also first finetuned on the full dataset with the same hyperparameters: 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout parameter.
-After that, unlearning techniques are applied to the model on the forget data using the following hyperparameters: 5 epochs of unlearning, batch size of 1, 2 gradient accumulation steps, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient.
+The "Gold" model is trained on the retain data only with 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$ and 0 dropout parameter.
+Unlearned models are also first finetuned on the full dataset with the same hyperparameters: 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout parameter.
+After that, unlearning techniques are applied to the model on the forget data using the following hyperparameters: 5 epochs of unlearning, batch size of 1, 2 gradient accumulation steps, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient.
 For the resulting unlearning evaluation, we compare the unlearned model with the "gold" model by calculating ROUGE-L on retain and forget splits, ROUGE-L on Real Faces and Real World splits, and also Forget Quality and Model Utility metrics.
-Appendix G U-MIA and U-LIRA
+Appendix I U-MIA and U-LIRA
 In this section, we provide details on evaluating unlearning methods using Unlearning Membership Inference Attack (U-MIA) algorithms.

### v4

@@
 We conduct a comprehensive analysis of 11 MU methods (e.g., SCRUB, gradient ascent, DPO) across four evaluation sets, demonstrating that jointly unlearning both modalities outperforms single-modality approaches.
+The dataset is available at https://huggingface.co/datasets/therem/CLEAR.
 CLEAR: Character Unlearning in Textual and Visual Modalities
-Alexey Dontsov1,6, Dmitrii Korzh1,3, Alexey Zhavoronkin2,4, Boris Mikheev3, Denis Bobkov1,6, Aibek Alanov1,6, Oleg Y.
-Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,6 1AIRI 2MIPT 3Skoltech 4Sber 5MTUCI 6HSE University Correspondence: dontsov@airi.net; tutubalina@airi.net
+Alexey Dontsov2,1, Dmitrii Korzh1,3, Alexey Zhavoronkin4,6, Boris Mikheev3, Denis Bobkov1,2, Aibek Alanov1,2, Oleg Y.
+Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,4,7 1AIRI 2HSE University 3Skoltech 4Sber AI 5MTUCI 6MIPT 7ISP RAS Research Center for Trusted AI Correspondence: dontsov@airi.net; tutubalina@airi.net
 1 Introduction
@@
 Finally, we compile a leaderboard of unlearning methods based on these evaluations.
-While MU has emerged as a promising alternative to costly retraining—enabling targeted removal of problematic data—existing methods focus almost exclusively on single-modality models.
+MU has emerged as a promising alternative to costly retraining, existing methods focus almost exclusively on single-modality models.
 Recent work has studied unlearning in LLMs Yao et al.
@@
 Recently, Chakraborty et al.
-(2024) pioneers the investigation of unlearning configurations in visual-language models (VLMs) to mitigate cross-modal safety risks, its experimental framework inherits a critical limitation: the datasets used (e.g., PKU-SafeRLHF Ji et al.
+(2024) pioneers the investigation of unlearning configurations in visual-language models (VLMs) to mitigate cross-modal safety risks, its experimental framework inherits a critical limitation: the datasets used in this study (e.g., PKU-SafeRLHF Ji et al.
 (2023), JailBreakV-28K Luo et al.
@@
 It includes forget/retain sets and real-world tasks (e.g., celebrity recognition) to evaluate cross-modal capability preservation.
-  • We comprehensively evaluate 11 recently proposed MU methods on our dataset and show that leading unimodal MU methods struggle in multimodal setups.
+  • We comprehensively evaluate 11 recently proposed MU methods on our dataset and show that the leading unimodal MU methods struggle in multimodal setups.
   • We establish leaderboards for textual, visual, and multimodal unlearning.
-We make all the data publicly available111at https://huggingface.co/datasets/fd28/CLEAR.
+We make all data publicly available111 https://huggingface.co/datasets/therem/CLEAR.
 2 Related Work
@@
 Maini et al.
-(2024) propose a benchmark named TOFU for textual LLM unlearning, consisting of 200 fictitious author profiles defined by attributes like name, birthplace, parent’s names, occupation, and written books, totalling 4,000 question-answer pairs (20 per author).
+(2024) propose a benchmark named TOFU for textual LLM unlearning, consisting of 200 fictitious author profiles defined by attributes like name, birthplace, parent’s names, occupation, and written books, totaling 4,000 question-answer pairs (20 per author).
 WMDP Li et al.
@@
 Our research investigates the multimodal aspects of unlearning.
+Figure 2: The overview of our multimodal dataset.
+The dataset consists of four sets: retain set, forget set, real faces (knowledge of related concepts such as faces), and real world (to evaluate general visual capabilities).
 Multimodality.
@@
 It uses CLIP Radford et al.
-(2021) to detect hallucinations based on MSCOCO-calibrated thresholds, eliminating manual labelling.
+(2021) to detect hallucinations based on MSCOCO-calibrated thresholds, eliminating manual labeling.
 The method applies three losses: negative loss to forget hallucinations, positive loss to reinforce correct representations, and sentence loss to preserve fluency.
@@
 Additionally, we utilize a holdout set $D_{H}$ such that $D_{H}\cap D=\emptyset$ to establish a reference for the model’s desired behaviour on $D_{F}$ after the unlearning process.
-In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behaviour on data that was not involved in the training process.
+In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behavior on data that was not involved in the training process.
 Such forgetting procedure is performed by updating the model $f_{\theta}$ with a particular unlearning method, which results in a new unlearned model $f_{\hat{\theta}}$ with parameters $\hat{\theta}$.
@@
 Yet, not all knowledge tied to $\mathcal{I}$ necessarily is required to be forgotten in this process.
-In multi-modal unlearning, compared to unimodal unlearning, each sample contains multiple modalities, e.g., text and image pairs. Let:
+In multi-modal unlearning, compared to unimodal unlearning, each sample contains multiple modalities, e.g., text and image pairs, and at least two modalities are processed for the MMU. Let:
 $D\;=\;\bigl{\{}\,\bigl{(}x_{i}^{\text{(1)}},\dots,x_{i}^{\text{(m)}},y_{i}\bigr{)}\bigr{\}}_{i=1}^{N},$ (4)
@@
 The same forget and retain goals hold.
+Note that multimodal data allows unimodal unlearning (for example, one can mask tokens of omitted modalities), but multimodal unlearning requires processing of several modalities.
+From the practical point of view, both MU and MMU depend on optimization loss functions, which at the same time depend on the inputs, either unimodal or multimodal.
+Some methods, originally proposed for the MU, can naturally be applied for the MMU; however, some MU methods, for example RMU or $\text{SCRUB}_{\text{bio}}$ are non-trivial to be adapted for the MMU.
 3.2 Methods
@@
 (2024).
-The objective, $L_{\text{IDK}}=L(D_{R},\theta)+L(D_{F}^{\text{idk}},\theta)$, ensures that the model retains performance on $D_{R}$ while aligning predictions on $D_{F}$ with uncertainty-based responses.
+The objective $L_{idk}$ ensures that the model retains performance on $D_{R}$ while aligning predictions on $D_{F}$ with uncertainty-based responses:
+$L_{idk}=L(D_{R},\theta)+L(D_{F}^{idk},\theta)$
 LLMU
 was introduced in early LLM unlearning research Yao et al. (2024c), optimizes the loss:
-$L_{\text{LLMU}}=-L(D_{F},\theta)+L(D_{F}^{\text{idk}},\theta)\\ +\sum_{x,y\in D_{R}}\operatorname{KL}(p_{\theta}(y|x)||p_{\hat{\theta}}(y|x)).$
+$L_{\text{LLMU}}=-L(D_{F},\theta)+L(D_{F}^{\text{idk}},\theta)\\ +\sum_{x,y\in D_{R}}\operatorname{KL}(p_{\theta}(y|x)||p_{\hat{\theta}}(y|x))$
 Here, $\theta$ and $\hat{\theta}$ are model’s parameters before and after unlearning, the first term promotes unlearning by maximizing loss on $D_{F}$, while the second reinforces forgetting using ”I don’t know” labels instead of original targets.
@@
 The student is optimized to match the teacher on the retain set $D_{R}$ while deviating on the forget set $D_{F}$.
-The loss function combines KL-divergence for retention ($L_{R}$), enforced divergence for unlearning ($L_{F}$), and task loss ($L_{\text{task}}$), yielding $L_{\text{SCRUB}}=L_{R}-L_{F}+L_{\text{task}}$.
+The loss function combines KL-divergence for retention ($L_{R}$), enforced divergence for unlearning ($L_{F}$), and task loss ($L_{\text{task}}$):
+$d(x,w^{s})=\operatorname{KL}(p(f(x;w^{o}))||p(f(x;w^{s}))),$
+$L_{R}=\frac{\alpha}{|D_{R}|}\sum_{x_{r}\in D_{R}}d(x_{r},w^{s}),$
+$L_{F}=\frac{1}{|D_{F}|}\sum_{x_{f}\in D_{F}}d(x_{f},w^{s}),$
+$L_{\text{task}}=\frac{\gamma}{|D_{R}|}\sum_{x_{r}\in D_{R}}l(x_{r},y_{r}),$
+$L_{\text{SCRUB}}=L_{R}-L_{F}+L_{\text{task}}$
+where $f(x;w^{o})$ is the original teacher model with weights $w^{o}$, which are kept unchanged, $f(x;w^{s})$ is the unlearned student model with parameters $w^{s}$, which are optimized.
 Preference Optimization (DPO)
@@
 (2023) for MU and MMMU.
-The model is trained to reduce reliance on undesired information through a loss function combining task performance retention ($L_{\text{task}}$) and a DPO-based loss ($L_{\text{DPO}}$), which penalizes deviations from a reference model $\pi_{ref}$ fine-tuned on $D_{F}^{\text{idk}}$ with ”I don’t know” labels as in IDK-tuning.
+The model is trained to reduce reliance on undesired information through a loss function combining task performance retention ($L_{\text{task}}$) and a DPO-based loss ($L_{\text{DPO}}$), which penalizes deviations from a reference model $\pi_{ref}$ fine-tuned on $D_{F}^{\text{idk}}$ with “I don’t know” labels as in IDK-tuning.
 To sum up, we chose recently published MU methods for their easy adaptation to new modalities, needing only input data changes (text, images, or both) while maintaining core functionality.
 4 CLEAR
-Figure 2: Distributions of the attributes of the author’s faces. We show that CLEAR is balanced and representative regarding age, gender, and ethnicity.
+Figure 3: Examples of generated images showcasing a distinct individual from our dataset.
 In this section, we describe a new benchmark CLEAR designed for character unlearning.
@@
 The final distribution of face and author characteristics is shown in Fig.
-2.
+4.
 After matching each author with a face, we used the diffusion model Li et al.
@@
 C).
-Figure 3: The overview of our dataset.
 We perform a simple reality check to ensure the quality of generated faces.
-We use the CLIP ViT-L/14 model, usually considered a visual encoder for most of the VLMs, to get embeddings of these three image sets – our faces, CelebA Liu et al.
+We use the CLIP ViT-L/14 model, which is usually considered a visual encoder for most VLMs, to get embeddings of these three image sets – our faces, CelebA Liu et al.
 (2015) and WebFace Yi et al.
@@
 This shows that the distance between our faces and the real-world faces is comparable to the distance between two real-world face datasets.
-Besides the author’s face, the diffusion model needs a textual prompt to produce an image.
+In addition to the author’s face, the diffusion model needs a textual prompt to produce an image.
 We ask GPT-4 to generate these prompts from a question-answer pair from TOFU about an author.
-We generate 8 images for each prompt, evaluate them using an ensemble of fake-detection models, and select the most realistic one.
+We generate $8$ images for each prompt, evaluate them using an ensemble of fake detection models, and select the most realistic one.
 Additionally, GPT-4o generates captions for each (image, visual prompt) pair, which are then included in the dataset to form pairs (image, caption).
@@
 4.2 Splits
+Figure 4: Distributions of the attributes of the author’s faces. We show that CLEAR is balanced and representative regarding age, gender, and ethnicity.
 We utilize four splits (sets) to evaluate MU (see Fig. 2 for a sample from these splits):
@@
 To ensure that the model’s overall visual capabilities remain intact throughout the unlearning process, we evaluate its performance on the Visual Question Answering (VQA) task using samples from x.ai (2024).
-Figure 4: Examples of generated images showcasing a distinct individual from our dataset.
 5 Experimental Setup and Evaluation
@@
 (2024a); Xing et al.
-(2024), this evaluation setup ensures that we capture the effectiveness of unlearning algorithms while examining both retention and forgetfulness of information within models.
+(2024), this evaluation setup ensures that we capture the effectiveness of unlearning algorithms while examining both the retention and forgetfulness of information within the models.
 ROUGE-L.
@@
 We set up the Real, Retain, and Forget metrics as a harmonic mean of the ROUGE, the Probability score, and the Truth Ratio computed on corresponding dataset splits.
-Forget Quality calculates the "distance" of the unlearned model to the gold model, which is a proxy metric for the quality of exact unlearning.
+Forget Quality calculates the “distance” of the unlearned model to the gold model, which is a proxy metric for the quality of exact unlearning.
 Following Maini et al.
 (2024), we take the Truth Ratios distribution of both models.
-But, instead of the p-value of the Kolmogorov-Smirnov test, we calculate the Jensen-Shannonn distance between these distributions.
-The latter metric better captures the differences between models, which we additionally check and describe in the Appx.
+However, instead of the p-value of the Kolmogorov-Smirnov test, we calculate the Jensen-Shannonn distance between these distributions.
+The latter metric better captures the differences between models, which we additionally check and describe in Appx.
 E.
 To maintain the higher - the better convention, we subtract the distance from 1.
-Method Text-only Visual-only Multimodal
-(Forget/Retain) (Forget/Retain) (Forget/Retain)
-LLMU 0.01/0.03 85.2/88.9 0.25/0.51
-DPO 0.42/0.26 50.2/81.4 0.22/0.48
-SCRUB 0.42/0.26 42.59/99.4 0.36/0.52
-IDK 0.24/0.26 N/A 0.33/0.51
-RMU 0.59/0.26 67.9/99.0 0.00/0.00
-Retain FT 0.42/0.26 100.0/100.0 0.37/0.51
-Performance correlation with multimodal:
-Spearman’s $\rho$ 0.705 (p=0.118) 0.205 (p=0.741) 1.00
-Table 2: Transferability analysis across domains.
-We report forget (F) and retain (R) metrics for each method.
-Lower F and higher R are better.
-N/A indicates the method was not applicable.
-Correlation shows Spearman’s rank correlation between single-domain and multimodal performance, with p-values in parentheses.
 5.2 Implementation
@@
 For comparison, we also present the metrics of the “gold” model.
-Experiment results and metrics are shown in Tab.
+Experimental results and metrics are shown in Tab.
 1, with details provided in Appx.
 H.
+In addition, we perform experiments with the QwenVL2 model (Appx.
+I).
 6 Results
+In the following sections, we describe the results of MU methods on our dataset. We seek to answer the following research questions:
+  1. RQ1:
+    Does an unlearning method’s performance on a single domain transfer directly to the performance in a multimodal setting?
+Should we study multimodal unlearning at all if we can easily predict its performance from single-domain experiments?
+  2. RQ2:
+    Is unlearning only one modality (textual or visual) enough in a multimodal setup?
+How does the effectiveness vary depending on modality?
+How do different unlearning methods compare in their effectiveness for unlearning specific modalities?
+  3. RQ3:
+    In the context of multimodal unlearning, what methods perform the best?
+6.1 Transferability from Single Domain
+To investigate RQ1, we analyze how well the performance of unlearning methods in single modalities predicts their effectiveness in multimodal settings.
+For the textual domain, we use the TOFU benchmark; for the visual, we use a standard U-MIA approach to the data, consisting of the faces from our dataset; the details for the pipelines and full results are provided in Appx.
+F and G.
+For multimodal unlearning, we use our benchmark.
+The correlations between single-domain and multimodal (MM) rankings are relatively weak.
+We rank methods according to their forget metric performance in each domain and calculate Spearman’s rank correlation coefficient between single-domain and multimodal rankings.
+The correlation $\rho=0.7$ for text-MM and $\rho=0.2$ for visual-MM, indicating limited transferability (see Tab.
+2).
+We observe significant discrepancies between single-domain and multimodal performance.
+For example, LLMU achieves a retain metric of 0.51 in multimodal but degrades to 0.03 in the textual setting while maintaining good forget scores (0.25 vs 0.01).
+Similar patterns emerge for other methods, suggesting that single-domain evaluation is insufficient.
+Methods that perform well in single domains can fail catastrophically in multimodal settings.
+For instance, RMU achieves good forget-retain balance in text-only (0.26/0.59) but completely fails on the retain set in multimodal setup (0.00/0.00), highlighting the unique challenges of multimodal unlearning.
+Takeaway 1: Single-domain performance is a poor predictor of multimodal unlearning success, with relatively low-rank correlations ($\rho=0.7$ and $\rho=0.2$) and distinct failure modes.
+This emphasizes the need for dedicated multimodal evaluation frameworks and potentially new methods designed specifically for multimodal MU.
+Method Text-only Visual-only Multimodal
+(Forget/Retain) (Forget/Retain) (Forget/Retain)
+LLMU 0.01/0.03 85.2/88.9 0.25/0.51
+DPO 0.42/0.26 50.2/81.4 0.22/0.48
+SCRUB 0.42/0.26 42.59/99.4 0.36/0.52
+IDK 0.24/0.26 N/A 0.33/0.51
+RMU 0.59/0.26 67.9/99.0 0.00/0.00
+Retain FT 0.42/0.26 100.0/100.0 0.37/0.51
+Performance correlation with multimodal:
+Spearman’s $\rho$ 0.705 0.205 1.00
+Table 2: Transferability analysis across domains.
+We report forget (F) and retain (R) metrics for each method.
+Lower F and higher R are better.
+N/A indicates the method was not applicable.
+Correlation shows Spearman’s rank correlation between single-domain and multimodal performance, with p-values in parentheses.
+6.2 Impact of Modality Selection on Unlearning
+To explore RQ2, we examine how the choice of unlearning modality impacts performance.
+We conduct experiments with three variants for each method: text-only, visual-only, and both modalities.
+Results in Tab.
+3 show distinct patterns across methods.
+Text-only Unlearning.
+Text-only approaches show mixed results.
+While some methods, like NPO, achieve good retain metrics (0.51) with moderate forget scores (0.29), others struggle significantly.
+RMU and GD completely fail on retain (0.01 and 0.00).
+KL shows middling performance with retain at 0.32 and forget at 0.23.
+This inconsistency suggests that text-only unlearning may disrupt cross-modal representations unpredictably.
+Visual-only Unlearning.
+Visual-only unlearning often achieves a better balance.
+DPO shows promising results with a forget metric of 0.23 while maintaining a 0.49 retain score.
+LLMU and SCRUB demonstrate similar patterns (forget: 0.37, 0.39; retain: 0.50, 0.49, respectively).
+However, some methods like RMU, GD, GA, and KL completely fail on retain metrics (0.00), indicating visual-only approaches are not universally successful.
+Multimodal Unlearning.
+Joint modality unlearning shows the most promising results for several methods.
+IDK improves its forget metric from 0.39 (text) and 0.35 (visual) to 0.33 (both) while maintaining stable retain performance (0.46).
+NPO shows strong real-world performance (0.49) but struggles with retain metrics when using both modalities.
+SCRUB demonstrates remarkable consistency across configurations (retain:  0.50, forget:  0.37), suggesting some methods are more robust to modality selection.
 Method Modality Real $\uparrow$ metric Retain $\uparrow$ metric Forget $\downarrow$ metric Forget $\uparrow$ Quality
@@
 Gold - a model trained only on retain.
-In the following sections, we describe the results of MU methods on our dataset. We seek to answer the following research questions:
-  1. RQ1:
-    Does an unlearning method’s performance on a single domain transfer directly to the performance in a multimodal setting?
-Should we study multimodal unlearning at all if we can easily predict its performance from single-domain experiments?
-  2. RQ2:
-    Is unlearning only one modality (textual or visual) enough in a multimodal setup?
-How does the effectiveness vary depending on modality?
-How do different unlearning methods compare in their effectiveness for unlearning specific modalities?
-  3. RQ3:
-    In the context of multimodal unlearning, what methods perform the best?
-6.1 Transferability from Single Domain
-To investigate RQ1, we analyse how well the performance of unlearning methods in single modalities predicts their effectiveness in multimodal settings.
-For the textual domain, we use the TOFU benchmark; for the visual, we use a standard U-MIA approach to the data, consisting of the faces from our dataset; the details for the pipelines and full results are provided in Appx.
-F and G.
-For multimodal unlearning, we use our benchmark.
-The correlations between single-domain and multimodal (MM) rankings are relatively weak.
-We rank methods according to their forget metric performance in each domain and calculate Spearman’s rank correlation coefficient between single-domain and multimodal rankings.
-The correlation $\rho=0.7$ for text-MM and $\rho=0.2$ for visual-MM, indicating limited transferability (see Tab.
-2).
-We observe significant discrepancies between single-domain and multimodal performance.
-For example, LLMU achieves a retain metric of 0.51 in multimodal but degrades to 0.03 in the textual setting while maintaining good forget scores (0.25 vs 0.01).
-Similar patterns emerge for other methods, suggesting that single-domain evaluation is insufficient.
-Methods that perform well in single domains can fail catastrophically in multimodal settings.
-For instance, RMU achieves good forget-retain balance in text-only (0.26/0.59) but completely fails on the retain set in multimodal setup (0.00/0.00), highlighting the unique challenges of multimodal unlearning.
-Takeaway 1: Single-domain performance is a poor predictor of multimodal unlearning success, with relatively low-rank correlations ($\rho=0.7$ and $\rho=0.2$) and distinct failure modes.
-This emphasizes the need for dedicated multimodal evaluation frameworks and potentially new methods designed specifically for multimodal MU.
-6.2 Impact of Modality Selection on Unlearning
-To explore RQ2, we examine how the choice of unlearning modality impacts performance.
-We conduct experiments with three variants for each method: text-only, visual-only, and both modalities.
-Results in Tab.
-3 show distinct patterns across methods.
-Text-only Unlearning.
-Text-only approaches show mixed results.
-While some methods, like NPO, achieve good retain metrics (0.51) with moderate forget scores (0.29), others struggle significantly.
-RMU and GD completely fail on retain (0.01 and 0.00).
-KL shows middling performance with retain at 0.32 and forget at 0.23.
-This inconsistency suggests that text-only unlearning may disrupt cross-modal representations unpredictably.
-Visual-only Unlearning.
-Visual-only unlearning often achieves a better balance.
-DPO shows promising results with a forget metric of 0.23 while maintaining a 0.49 retain score.
-LLMU and SCRUB demonstrate similar patterns (forget: 0.37, 0.39; retain: 0.50, 0.49, respectively).
-However, some methods like RMU, GD, GA, and KL completely fail on retain metrics (0.00), indicating visual-only approaches are not universally successful.
-Multimodal Unlearning.
-Joint modality unlearning shows the most promising results for several methods.
-IDK improves its forget metric from 0.39 (text) and 0.35 (visual) to 0.33 (both) while maintaining stable retain performance (0.46).
-NPO shows strong real-world performance (0.49) but struggles with retain metrics when using both modalities.
-SCRUB demonstrates remarkable consistency across configurations (retain:  0.50, forget:  0.37), suggesting some methods are more robust to modality selection.
 Takeaway 2: While visual-only unlearning often outperforms text-only approaches, the effectiveness varies significantly by method.
@@
 Consequently, certain sections of our publication may be identified as AI-generated, AI-edited, or a combination of human and AI contributions.
+Acknowledgements
+The work was supported by a Research Center for Trusted Artificial Intelligence the Ivannikov Institute for System Programming of the Russian Academy of Sciences.
 Appendix A Rationale for Synthetic Face Selection
@@
     Teacher-Student (SCRUB) Kurmanji et al.
-(2023) The main idea of this method is to train a student model, which is taken as a desired unlearned model from the original one, such that it will "disobey" the teacher original model on the forget set.
+(2023) The main idea of this method is to train a student model, which is taken as a desired unlearned model from the original one, such that it will “disobey” the teacher original model on the forget set.
 The resulting loss of student model in this method is constructed as follows:
@@
 Therefore, we do not consider these methods to be successful.
+Method Real Retain Forget Forget Quality
+Gold 0.27 0.66 0.05 1.00
+Original 0.27 0.65 0.40 0.97
+GD 0.20 0.15 0.05 0.36
+GA 0.17 0.10 0.05 0.89
+IDK 0.04 0.65 0.40 0.98
+KL 0.18 0.04 0.05 0.87
+LLMU 0.07 0.63 0.33 0.99
+NPO 0.16 0.20 0.13 0.97
+Retain FT 0.25 0.66 0.32 0.98
+SCRUB 0.25 0.65 0.39 0.98
+Table 6: Comparison of unlearning methods on the QwenVL2 model.
 Appendix D A sample of dataset
-Our dataset consists of 200 fictitious authors, each with 15-20 visual and 20 textual questions. We add an example of data for a single person in the Table 6.
+Our dataset consists of 200 fictitious authors, each with 15-20 visual and 20 textual questions. We add an example of data for a single person in Tab. 7.
 Image Caption
@@
 Such experimental settings and hyperparameters are the same for both Llama2-7B and Mistral architectures.
-To assess the unlearning quality, we compare the obtained unlearned model with the "gold" one and calculate ROUGE-L on retain and forget parts, Forget Quality and Model Utility metrics.
-The full results are available in table 4.
+To assess the unlearning quality, we compare the obtained unlearned model with the “gold” one and calculate ROUGE-L on retain and forget parts, Forget Quality and Model Utility metrics.
+Full results are available in Tab.
+4.
 Appendix G Visual-only Unlearning
@@
 For each of these 128 models, we run the forgetting algorithm on the forget subset of this particular model.
-From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see section I) on which the quality of the forgetting algorithms will be tested.
+From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see Appx.
+J) on which the quality of the forgetting algorithms will be tested.
 Each of the 64 target models forgets a sample $\mathcal{D}f$ of 20 personalities.
 Additionally, for each target model, we form a holdout set $D_{H}$ by selecting 20 personalities that were not used in the training of this model.
-The full results are available at table 5
+The full results are available in Table 5.
 In our experiments, we employ U-LIRA with 64 shadow models, with half representing the in-distribution and the other half representing the out-distribution for each target example.
@@
 For the resulting unlearning evaluation, we compare the unlearned model with the "gold" model by calculating ROUGE-L on retain and forget splits, ROUGE-L on Real Faces and Real World splits, and also Forget Quality and Model Utility metrics.
-Appendix I U-MIA and U-LIRA
+Appendix I Multimodal unlearning on QwenVL series
+In addition to our experiments on the LLAVA series, we provide results for the QwenVL2-2B model in the table 6. We use same hyperparameters as stated in H, except that we do not use the LoRA adapters.
+Appendix J U-MIA and U-LIRA
 In this section, we provide details on evaluating unlearning methods using Unlearning Membership Inference Attack (U-MIA) algorithms.